### v1

# Stationary solutions to stochastic 3D Euler equations in Hölder space

Abstract.
We establish the existence of infinitely many global and stationary solutions in $C(\mathbb{R};C^{\vartheta})$ space for some $\vartheta>0$ to the three dimensional Euler equations driven by an additive noise. The result is based on a new stochastic version of the convex integration method, incorporating the stochastic convex integration method developed in [HZZ22b] and pathwise estimates to derive uniform moment estimates independent of time.
Key words and phrases:
stochastic Euler equations, stationary solutions, Hölder space, convex integration.
2010 Mathematics Subject Classification:
60H15; 35R60; 35Q30
Research supported by National Key R&D Program of China (No. 2022YFA1006300) and the NSFC (No. 12271030). The financial support by the DFG through the CRC 1283 ”Taming uncertainty and profiting from randomness and low regularity in analysis, stochastics and their applications” is greatly acknowledged.
1. Introduction
In this paper, we are considered with stochastic Euler equations governing the time evolution of the velocity $u$ of an inviscid fluid on the three dimensional torus $\mathbb{T}^{3}=\mathbb{R}^{3}/(2\pi\mathbb{Z})^{3}$. The system reads as
${\mathord{{\rm d}}}u+\mathord{{\rm div}}(u\otimes u)\,{\mathord{{\rm d}}}t+\nabla P\,{\mathord{{\rm d}}}t$ $={\mathord{{\rm d}}}B,$ (1.1)
$\mathord{{\rm div}}u$ $=0,$
where $P$ stands for the corresponding pressure. The right hand side represents a random external force acting on the fluid, which is given by a $GG^{*}$-Wiener process $B$ defined on some probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P})$ with $GG^{*}$ being trace class operator. In early works, stochastic Euler equations have been widely studied in both two and three dimensions, see e.g.[BFM16], [BP01], [CFH19], [GHV14], where local-in-time solutions for (1.1) have been constructed. Our goal is to investigate the existence of global and stationary solutions with Hölder continuity to the above system (1.1). To achieve this, we employ the method of convex integration, a technique that has gained prominence in fluid dynamics over the past decades. This method was initially introduced following [DLS10, DLS12, DLS13] by De Lellis and Székelyhidi Jr. In these works, the method of convex integration was used in order to prove non-uniqueness of weak solutions with Hölder continuity to Euler equations. Subsequently, this method has led to several breakthroughs in the fluid dynamics, including proof of Onsager’s conjecture for the incompressible Euler equations [Ise18, BDLSV19]. We refer to the excellent review article by Buckmaster and Vicol [BV19] for a gentle introduction and further references. In recent years, convex integration has already been applied in a stochastic setting, namely, to the isentropic Euler system by Breit, Feireisl and Hofmanová [BFH20] and to the full Euler system by Chiodaroli, Feireisl and Flandoli [CFF19] with linear multiplicative noise. Futhermore, ill/well-posedness of dissipative martingale solutions to stochastic 3D Euler equations was studied by Hofmanová, Zhu and Zhu [HZZ22a], along with the existence and non-uniqueness of strong Markov solutions. Later, Hölder continuous, global-in-time probabilistically strong solutions to 3D Euler equations perturbed by Stratonovich transport noise was constructed in [HLP22]. In the mostly recent work [HZZ22b], existence and non-uniqueness of stationary solutions in $L^{2}$ to the stochastic 3D Euler equations was established, using a new stochastic convex integration method. However, the regularity of the global and stationary solutions to the stochastic 3D Euler equations with additive noise, as constructed in the previous work, cannot achieve Hölder continuity. Inspired by these previous works and the progress in the deterministic case, it is natural to investigate the existence and non-uniqueness of global and stationary Hölder continuous solutions to the Euler equations in the stochastic setting (1.1). Another motivation for our research is linked to the Onsager’s conjecture. It is constructed in [Ise18, BDLSV19] that solutions in $C^{\frac{1}{3}-}$ do not conserve energy to the deterministic Euler equations. It is natural to ask whether the regularity of the global, nonunique solutions in the stochastic setting can stay $C^{\frac{1}{3}-}$ a.s.. To summarize the above discussion, it is natural to investigate the validity of the following claims.
  1. (i)
    Existence and non-uniqueness of global $\vartheta-$Hölder continuous solutions to the stochastic Euler equations (1.1), for some $\vartheta>0$.
  2. (ii)
    Existence and non-uniqueness of stationary solutions to the stochastic Euler equations (1.1).
1.1. Main results
Our aim is to provide some answers to the above problems (i) and (ii) in the physically relevant context of the stochastic Euler equations on $\mathbb{T}^{3}$ driven by additive stochastic noise. As the main tool within this study, we apply a stochastic version of the convex integration method. Unlike the previous works using convex integration for the Navier–Stokes and Euler equations in the stochastic setting [HZZ19, HZZ22a, HZZ23a, HZZ23b], we are now inspired by [CDZ22] and [HZZ22b] where they overcome the limitation originating from stopping times, previously used to control the noise terms in the iteration. More precisely, we no longer work with stopping times and instead we include expectations in the inductive iteration and provide uniform moment estimates. However, the construction of Hölder continuous Euler flows in the deteministic setting, as demonstrated in [BV19], requires to introduce transport equations, and pathwise positive lower bounds for the solutions are used to improve regularity. As a result, we also introduce some pathwise estimates during iteration for convex integration. To this end, we use cutoff method to control the growth of noise pathwisely. Compared to only using stochastic convex integration, the combination of pathwise estimates with it can provide better Hölder regularity (see Remark 1.3 below). This approach also allows us not to control all the moments bounds during the iteration. We hope this cutoff method can be applied to enhance regularity of the solutions in the future work. As we are interested in the long time behavior of solutions and particularly in the construction of statistically stationary solutions, we work with entire solutions solving the equations for all times $t\in{\mathbb{R}}$. Accordingly, the norms in the convex integration scheme must be chosen appropriately in order to provide the desired global-in-time estimates. This is achieved through bounds of the form
$\sup_{t\in\mathbb{R}}\mathbf{E}\left[\sup_{t\leqslant s\leqslant t+1}\|u(s)\|_{C^{\vartheta}}^{p}\right]<\infty,$
with $p\geqslant 1$ and small $\vartheta>0$. Such bounds provide uniform moment estimates locally in $C(\mathbb{R};C^{\vartheta})$ and guarantee the convergence of the corresponding ergodic averages. This leads to the existence of stationary solutions. Within this study, we focus on analytically weak solutions which satisfy the equations in the following sense.
Definition 1.1.
We say that $((\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P}),u,B)$ is an analytically weak solution to the Euler system (1.1) provided
  1. (1)
    $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P})$ is a stochastic basis with a complete right-continuous filtration;
  2. (2)
    $B$ is an $\mathbb{R}^{3}$-valued, spatial mean and divergence free, two-sided $GG^{*}$-Wiener process with respect to the filtration $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$;
  3. (3)
    the velocity $u\in C({\mathbb{R}}\times\mathbb{T}^{3})$ $\mathbf{P}$-a.s. and is $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted;
  4. (4)
    for every $-\infty<s\leqslant t<\infty$ it holds $\mathbf{P}$-a.s.
    $\langle u(t),\psi\rangle+\int_{s}^{t}\langle\mathord{{\rm div}}(u\otimes u),\psi\rangle{\mathord{{\rm d}}}r\ =\langle u(s),\psi\rangle+\langle B(t)-B(s),\psi\rangle$
    for all $\psi\in C^{\infty}(\mathbb{T}^{3})$, $\mathord{{\rm div}}\psi=0$.
With this definition at hand, our first main result reads as follows, tackling the problem (i) above.
Theorem 1.2.
Suppose that $\mathrm{Tr}((-\Delta)^{\frac{3}{2}+\sigma}GG^{*})<\infty$ for some $\sigma>0$, then there exists an analytically weak solution to (1.1) in the sense of Definition 1.1 which belongs to $C(\mathbb{R},C^{\vartheta})$, for some $\vartheta\in\left(0,{\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot}7^{5}},\frac{1}{3{\cdot}7^{5}}\}}\right)$.
Remark 1.3.
(i) In Section 3, we prove this result by combining stochastic convex integration method with pathwise estimates to derive uniform moment estimates independent of time, and obtain Hölder continuous solutions, where $\vartheta$ can be chosen less than ${\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot}7^{5}},\frac{1}{3{\cdot}7^{5}}\}}$. While we also prove this result in a separable article (see [Lü24]), which applies stochastic convex integration method to the the construction of Hölder continuous Euler flows in [DLS13] to derive uniform moment estimates independent of time. The value of $\vartheta$ in this case decreases due to the higher moments required, and we may take $\vartheta<{\mathord{{\rm min}}}{\{\frac{\sigma}{73{\cdot}62^{3}},\frac{1}{7{\cdot}62^{3}}\}}$. (ii) It is natural to ask whether we could construct the solutions with better regulartiy to (1.1) by combining the gluing method in [BDLSV19] with our cutoff method. This will be studied in our future work.
The second result is the existence and non-uniqueness of stationary solutions in Hölder space. In the sequel, we understand stationarity in the sense of shift invariance of laws of solutions on the space of trajectories, see also [BFHM19, BFH20e, FFH21, HZZ22b]. More precisely, we define the joint trajectory space for the solution and the driving Brownian motion as
${\mathcal{T}}=C({\mathbb{R}};C^{\kappa})\times C({\mathbb{R}};C^{\kappa})$
for some $\kappa>0$ and let $S_{t}$, $t\in{\mathbb{R}}$, be shifts on trajectories given by
$S_{t}(u,B)(\cdot)=(u(\cdot+t),B(\cdot+t)-B(t)),\quad t\in{\mathbb{R}},\quad(u,B)\in{\mathcal{T}}.$
We note that the shift in the second component acts differently in order to guarantee that for a Brownian motion $B$ the shift $S_{t}B$ is again a Brownian motion.
Definition 1.4.
We say that $((\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P}),u,B)$ is a stationary solution to the stochastic Euler equations (1.1) provided it satisfies (1.1) in the sense of Definition 1.1 and its law is shift invariant, that is,
$\mathcal{L}[S_{t}(u,B)]=\mathcal{L}[u,B]\qquad\text{ for all }\quad t\in{\mathbb{R}}.$
With these definitions at hand, we are able to state our second main result, tackling the problem (ii) above. The result then reads as follows and is proved in Section 5.
Theorem 1.5.
Suppose that $\mathrm{Tr}((-\Delta)^{\frac{3}{2}+\sigma}GG^{*})<\infty$ for some $\sigma>0$, then there exist infinitely many stationary solutions to the stochastic Euler equations (1.1). Moreover, the solutions belong to $C({\mathbb{R}},C^{\vartheta})$ a.s. for some $\vartheta>0$, satisfying
$\sup_{t\in\mathbb{R}}\mathbf{E}\left[\sup_{t\leqslant s\leqslant t+1}\|u(s)\|_{C^{\vartheta}}\right]<\infty.$
Organization of the paper
In Section 2, we collect the basic notations used throughout the paper. Section 3 and Section 4 are the core of our proofs: the stochastic convex integration is developed and employed to construct entire non-unique analytically weak solutions. In Section 3, we present the main iteration Proposition 3.2 to prove our main Theorem 3.3. Section 4 is devoted to the construction of stochastic convex integration with pathwise estimates, namely, proof of iteration Proposition 3.2. Then the results in Section 3 are used in Section 5 together with a Krylov–Bogoliubov’s argument to obtain existence of non-unique stationary solutions to the stochastic Euler equations. In Appendix A, we recall the construction of Beltrami waves from [BV19] . In Appendix B, we give estimates on amplitude functions used in the convex integration construction. Finally, the estimates for transport equations are completed in Appendix C.
2. Notations
2.1. Function spaces
Throughout the paper, we employ the notation $a\lesssim b$ if there exists a constant $c>0$ such that $a\leqslant cb$. We let ${\mathbb{N}}_{0}:={\mathbb{N}}\cup\{0\}$. Given a Banach space $E$ with a norm $\|\cdot\|_{E}$ and $t\in{\mathbb{R}}$, we write $C_{t}E=C([t,t+1];E)$ for the space of continuous functions from $[t,t+1]$ to $E$, equipped with the supremum norm $\|f\|_{C_{t}E}=\sup_{s\in[t,t+1]}\|f(s)\|_{E}$. For $\alpha\in(0,1)$ we define $C^{\alpha}_{t}E$ as the space of $\alpha$-Hölder continuous functions from $[t,t+1]$ to $E$, endowed with the norm $\|f\|_{C^{\alpha}_{t}E}=\sup_{s,r\in[t,t+1],s\neq r}\frac{\|f(s)-f(r)\|_{E}}{|r-s|^{\alpha}}+\sup_{s\in[t,t+1]}\|f(s)\|_{E}$. For $\alpha\in(0,1)$, we also write $C^{\alpha}$ for functions in the space of $\alpha$-Hölder continuous functions from $\mathbb{T}^{3}$ to ${\mathbb{R}}^{3}$ endowed with the norm $\|f\|_{C^{\alpha}}=\sup_{x,y\in\mathbb{T}^{3},x\neq y}\frac{|f(x)-f(y)|}{|x-y|^{\alpha}}+\sup_{x\in\mathbb{T}^{3}}|f(x)|.$ We use $L^{p}$ to denote the set of standard $L^{p}$-integrable functions from $\mathbb{T}^{3}$ to $\mathbb{R}^{3}$. For $s>0$, $p>1$ we set $W^{s,p}:=\{f\in L^{p};\|(I-\Delta)^{{s}/{2}}f\|_{L^{p}}<\infty\}$ with the norm $\|f\|_{W^{s,p}}=\|(I-\Delta)^{{s}/{2}}f\|_{L^{p}}$. Set $L^{2}_{\sigma}=\{f\in L^{2};\int_{\mathbb{T}^{3}}f\,{\mathord{{\rm d}}}x=0,\mathord{{\rm div}}f=0\}$. For $s>0$, we define $H^{s}:=W^{s,2}\cap L^{2}_{\sigma}$. For $s<0$ we define $H^{s}$ to be the dual space of $H^{-s}$. For $t\in\mathbb{R}$ and a domain $D\subset\mathbb{R}$ we denote by $C^{N}_{t,x}$ and $C^{N}_{D,x}$, respectively, the space of $C^{N}$-functions on $[t,t+1]\times\mathbb{T}^{3}$ and on $D\times\mathbb{T}^{3}$, respectively, $N\in\mathbb{N}_{0}$. The spaces are equipped with the norms
$\|f\|_{C^{N}_{t,x}}=\sum_{\begin{subarray}{c}0\leqslant n+|\alpha|\leqslant N\\ n\in\mathbb{N}_{0},\alpha\in\mathbb{N}^{3}_{0}\end{subarray}}\|\partial_{t}^{n}D^{\alpha}f\|_{L^{\infty}_{[t,t+1]}L^{\infty}},\qquad\|f\|_{C^{N}_{D,x}}=\sum_{\begin{subarray}{c}0\leqslant n+|\alpha|\leqslant N\\ n\in\mathbb{N}_{0},\alpha\in\mathbb{N}^{3}_{0}\end{subarray}}\sup_{t\in D}\|\partial_{t}^{n}D^{\alpha}f\|_{L^{\infty}}.$
For a Polish space $H$ we denote by $\mathcal{B}(H)$ the $\sigma$-algebra of Borel sets in $H$. We also use $\mathring{\otimes}$ to denote the trace-free part of the tensor product. We recall the inverse divergence operator $\mathcal{R}$ from [DLS13, Section 4.5], which acts on vector fields $v$ with $\int_{\mathbb{T}^{3}}v{\mathord{{\rm d}}}x=0$ as
$(\mathcal{R}v)^{kl}=(\partial_{k}\Delta^{-1}v^{l}+\partial_{l}\Delta^{-1}v^{k})-\frac{1}{2}(\delta_{kl}+\partial_{k}\partial_{l}\Delta^{-1})\mathord{{\rm div}}\Delta^{-1}v,$
for $k,l\in\{1,2,3\}$. Then $\mathcal{R}v(x)$ is a symmetric trace-free matrix for each $x\in\mathbb{T}^{3}$, and $\mathcal{R}$ is a right inverse of the div operator, i.e. $\mathord{{\rm div}}(\mathcal{R}v)=v$. By [CL22, Theorem B.3] we have for $1\leqslant p\leqslant\infty$
$\|\mathcal{R}f\|_{L^{p}(\mathbb{T}^{3})}\lesssim\|f\|_{L^{p}(\mathbb{T}^{3})}.$ (2.1)
2.2. Probabilistic elements
For a given probability measure $\mathbf{P}$ we denote by $\mathbf{E}$ the expectation under $\mathbf{P}$. Regarding the driving noise, we assume that $B$ is $\mathbb{R}^{3}$-valued two-sided $GG^{*}$-Wiener process with zero spatial mean and zero divergence, defined on some probability space $(\Omega,\mathcal{F},\mathbf{P})$ and $G$ is a Hilbert–Schmidt operator from $U$ to $L_{\sigma}^{2}$ for some Hilbert space $U$. For $p\in[1,\infty)$, $\delta\in(0,\frac{1}{2})$ and some $\kappa>0$, we denote
$|\!|\!|u|\!|\!|^{p}_{C^{0},p}:=\sup_{t\in\mathbb{R}}\mathbf{E}\left[\sup_{s\in[t,t+1]}\|u(s)\|^{p}_{C^{0}}\right],\quad|\!|\!|u|\!|\!|^{p}_{C_{t}^{\frac{1}{2}-\delta}C^{\kappa},p}:=\sup_{t\in{\mathbb{R}}}\mathbf{E}\left[\|u\|^{p}_{C_{t}^{\frac{1}{2}-\delta}C^{\kappa}}\right].$
These norms define function spaces of random variables on $\Omega$ taking values in $C({\mathbb{R}}\times\mathbb{T}^{3})$ and $C^{\frac{1}{2}-\delta}({\mathbb{R}},C^{\kappa})$, respectively, with bounds in $L^{p}(\Omega;{C^{0}(I\times\mathbb{T}^{3}))}$ and $L^{p}(\Omega;{C^{\frac{1}{2}-\delta}(I,C^{\kappa})})$ for bounded interval $I\subset{\mathbb{R}}$. Furthermore, the bounds only depend on the length of the interval $I$, not on its location within ${\mathbb{R}}$. In the sequel, we simply say that $u$ has a uniform moment of order $p$ locally in $C^{0}({\mathbb{R}}\times\mathbb{T}^{3})$ provided $|\!|\!|u|\!|\!|^{p}_{C^{0},p}<\infty$. Similarly, we define the corresponding norms with $C^{0}$ replaced by $C^{\kappa}$, $C^{1}$, $L^{1}$ and $H^{\frac{3}{2}+\kappa}$.
3. Stochastic convex integration with pathwise estimates and results
In [CDZ22] and [HZZ22b], they presented a stochastic convex integration method constructing directly solutions on the whole time line $\mathbb{R}$ by introducing expectations to the iterative estimates in convex integration. In this section, we integrate their methods with pathwise estimates in the context of the stochastic Euler equations. Our goal is to construct a Hölder continuous Euler flow, mirroring the deterministic case as presented in [BV19], where the transport equations (see (4.7) below) are used to improve regularity. This method necessitates pathwise positive lower bounds for the solutions of the transport equations. As a result, moment bounds during the inductive iteration become insufficient. Hence, we introduce additional pathwise estimates in convex integration schemes. More precisely, our idea employs the cutoff method to control the growth of the noise, enabling the introduction of pathwise estimates during the inductive iteration. This is combined with moment bounds to derive desired uniform estimates. Compared to only using moment bounds in the inductive iteration in [CDZ22] and [HZZ22b], the advantage of this approach is that it only requires lower moments of the solutions, and avoids the requirements for higher moments as in [HZZ22b]. This can simplify the computations and enhance the regularity of the solutions. At the first step, we decompose stochastic Euler system (1.1) into two parts, one is linear and involves the noise, whereas the second one is a random PDE. More precisely, we consider the stochastic linear equation
${\mathord{{\rm d}}}z+z{\mathord{{\rm d}}}t$ $={\mathord{{\rm d}}}B,$ (3.1)
$\mathord{{\rm div}}z$ $=0,$
where $B$ is $\mathbb{R}^{3}$-valued two-sided trace-class $GG^{*}$-Wiener process with spatial zero mean and divergence free (see e.g. [LR15, Section 2.1]), and $v$ solves the nonlinear equation
$\partial_{t}v-z+\mathord{{\rm div}}((v+z)\otimes(v+z))+\nabla P$ $=0,$ (3.2)
$\mathord{{\rm div}}v$ $=0.$
Here, $z$ is divergence free by the assumptions on the noise $B$ and by $P$ we denote the pressure term associated to $v$. We include the linear zero order term $z$ on the left hand side of (3.1) in order to obtain a stabilization of the equation needed for the necessary global in time estimates. It will be seen in the course of the construction that the corresponding counter term in (3.2) will not cause any difficulties. Using the factorization method, it is standard to derive regularity of the stochastic convolution $z$ on a given stochastic basis $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P})$ with $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$ canonical filtration given in [LR15, Section 2.1]. In particular, the following result follows from [DPZ92, Theorem 5.16] together with the Kolmogorov continuity criterion.
Proposition 3.1.
Suppose that $\mathrm{Tr}((-\Delta)^{\frac{3}{2}+\kappa}GG^{*})<\infty$ for some $\kappa>0$. Then for any $\delta\in(0,\frac{1}{2})$, $p\geqslant 2$
$\sup_{t\in{\mathbb{R}}}\mathbf{E}\left[\|z\|^{p}_{C_{t}^{1/2-\delta}{C^{\kappa}}}\right]\lesssim\sup_{t\in{\mathbb{R}}}\mathbf{E}\left[\|z\|^{p}_{C_{t}^{1/2-\delta}H^{\frac{3}{2}+\kappa}}\right]\leqslant(p-1)^{p/2}L^{p},$ (3.3)
where $L\geqslant 1$ depends on $\mathrm{Tr}((-\Delta)^{3/2+\kappa}GG^{*})$, $\delta$ and is independent of $p$.
Proof.
We recall that the unique stationary solution to (3.1) has the explicit form $z(t)=\int_{-\infty}^{t}e^{s-t}{\mathord{{\rm d}}}B_{s}$. By [LR15, Section 2.1], the Wiener process $B$ can be written as $B=\sum_{k\in{\mathbb{N}}}\sqrt{c_{k}}\beta_{k}e_{k}$ for an orthonormal basis $\{e_{k}\}_{k\in{\mathbb{N}}}$ of $L^{2}_{\sigma}$ consisting of eigenvectors of $GG^{*}$ with corresponding eigenvalues $c_{k}$, and the coefficients satisfy $\sum_{k\in{\mathbb{N}}}c_{k}<\infty$, where $\{\beta_{k}\}_{k\in{\mathbb{N}}}$ are a sequence of mutually independent standard two-sided real-valued Brownian motions. Then it holds for $t\geqslant s$,
$\mathbf{E}\|z(t)-z(s)\|^{2}_{H^{\frac{3}{2}+\kappa}}$
$=\mathbf{E}\left\|(I-\Delta)^{\frac{1}{2}(\frac{3}{2}+\kappa)}\int_{s}^{t}e^{r-t}{\mathord{{\rm d}}}B_{r}\right\|_{L^{2}}^{2}+\mathbf{E}\left\|(I-\Delta)^{\frac{1}{2}(\frac{3}{2}+\kappa)}\int_{-\infty}^{s}(e^{r-t}-e^{r-s}){\mathord{{\rm d}}}B_{r}\right\|^{2}_{L^{2}}$
$=\sum_{k=1}^{\infty}c_{k}\int_{s}^{t}\|(I-\Delta)^{\frac{1}{2}(\frac{3}{2}+\kappa)}e_{k}\|_{L^{2}}^{2}e^{2(r-t)}{\mathord{{\rm d}}}r$
$\qquad+\sum_{k=1}^{\infty}c_{k}\int_{-\infty}^{s}\|(I-\Delta)^{\frac{1}{2}(\frac{3}{2}+\kappa)}e_{k}\|_{L^{2}}^{2}(e^{r-t}-e^{r-s})^{2}{\mathord{{\rm d}}}r$
$\leqslant\sum_{k=1}^{\infty}c_{k}\left(\|e_{k}\|_{L^{2}}^{2}+\|(-\Delta)^{\frac{1}{2}(\frac{3}{2}+\kappa)}e_{k}\|_{L^{2}}^{2}\right)\left(\int_{s}^{t}e^{2(r-t)}{\mathord{{\rm d}}}r+\int_{-\infty}^{s}(e^{r-t}-e^{r-s})^{2}{\mathord{{\rm d}}}r\right)$
$=(\mathrm{Tr}(GG^{*})+\mathrm{Tr}((-\Delta)^{3/2+\kappa}GG^{*}))\left(\int_{s}^{t}e^{2(r-t)}{\mathord{{\rm d}}}r+\int_{-\infty}^{s}(e^{r-t}-e^{r-s})^{2}{\mathord{{\rm d}}}r\right)$
$\leqslant(\mathrm{Tr}(GG^{*})+\mathrm{Tr}((-\Delta)^{3/2+\kappa}GG^{*}))(t-s).$
Using Gaussianity we have
$\mathbf{E}\|z(t)-z(s)\|^{p}_{H^{\frac{3}{2}+\kappa}}\leqslant(p-1)^{\frac{p}{2}}\Big{(}\mathbf{E}\|z(t)-z(s)\|_{H^{\frac{3}{2}+\kappa}}^{2}\Big{)}^{\frac{p}{2}}.$
The result then follows from Kolmogorov’s continuity criterion and fundamental Sobolev embedding $H^{\frac{3}{2}+\kappa}\hookrightarrow C^{\kappa}$, i.e. for any $\delta\in(0,\frac{1}{2})$
$\sup_{t\in{\mathbb{R}}}\mathbf{E}\left[\|z\|^{p}_{C_{t}^{1/2-\delta}{C^{\kappa}}}\right]\lesssim\sup_{t\in{\mathbb{R}}}\mathbf{E}\left[\|z\|^{p}_{C_{t}^{1/2-\delta}H^{\frac{3}{2}+\kappa}}\right]\leqslant(p-1)^{p/2}L^{p},$
where $L\geqslant 1$ depends on $\mathrm{Tr}((-\Delta)^{\frac{3}{2}+\kappa}GG^{*})$, $\delta$ and is independent of $p$. ∎
3.1. Stochastic convex integration set-up and iterative proposition
Let us now explain how the convex integration iteration is set up. We consider an increasing sequence $\{\lambda_{q}\}_{q\in\mathbb{N}_{0}}$ which diverges to $\infty$, and a bounded sequence $\{\delta_{q}\}_{q\in\mathbb{N}}$ which decreases to 00. We choose $a,b\in\mathbb{N}$ and $\beta\in(0,1)$ and let
$\lambda_{q}=a^{b^{q}},\quad\delta_{1}=3rL^{2},\quad\delta_{q}=\frac{1}{2}\lambda_{2}^{2\beta}\lambda_{q}^{-2\beta},\ q\geqslant 2.$
Here $\beta$ will be chosen sufficiently small, $a$ will be chosen sufficiently large, and more details on the choice of these parameters will be given below in the course of the construction. The iteration is indexed by a parameter $q\in\mathbb{N}_{0}$. At each step $q$, a pair $(v_{q},\mathring{R}_{q})$ is constructed to solve the following system
$\partial_{t}v_{q}-z_{q}+\mathord{{\rm div}}((v_{q}+z_{q})\otimes(v_{q}+z_{q}))+\nabla p_{q}$ $=\mathord{{\rm div}}\mathring{R}_{q},$ (3.4)
$\mathord{{\rm div}}v_{q}$ $=0.$
In the above we define
$z_{q}(t,x)=\chi_{q}\left(\|\tilde{z}_{q}(t)\|_{C^{1}_{x}}\right)\tilde{z}_{q}(t,x),$ (3.5)
where $\chi_{q}$ is a non-increasing smooth function such that
$\chi_{q}(x)=\begin{cases}1,&x\in[0,\lambda_{q+2}^{\gamma}],\\ 0,&x\in(\lambda_{q+3}^{\frac{1}{4}\gamma},\infty),\end{cases}$
with derivative bounded by $1$, which requires
$\frac{1}{\lambda_{q+3}^{\gamma/4}-\lambda_{q+2}^{\gamma}}\leqslant 1.$ (3.6)
And $\tilde{z}_{q}=\mathbb{P}_{\leqslant f(q)}z$ with $f(q)=\lambda_{q+1}^{\gamma/8}$. Here $\mathbb{P}_{\leqslant f(q)}$ is the Fourier multiplier operator, which projects a function onto its Fourier frequencies $\leqslant f(q)$ in absolute value, and $\gamma$ is another small parameter which will be fixed in Section 4.1. $\mathring{R}_{q}$ on the right hand of (3.4) is a $3\times 3$ matrix which is trace-free and we put the trace part into the pressure. Keeping (3.3) in mind, by the Sobolev embedding $\|f\|_{L^{\infty}}\lesssim\|f\|_{H^{\frac{3}{2}+\kappa}}$ for $\kappa>0$, we observe that for any $p\geqslant 2$ and $\delta\in(0,\frac{1}{2})$
$|\!|\!|z_{q}|\!|\!|_{C^{0},p}$ $\lesssim|\!|\!|\tilde{z}_{q}|\!|\!|_{H^{\frac{3}{2}+\kappa},p}\lesssim|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\kappa},p}\lesssim(p-1)^{\frac{1}{2}}L,$ (3.7)
$|\!|\!|z_{q}|\!|\!|_{C^{1},p}$ $\lesssim|\!|\!|\tilde{z}_{q}|\!|\!|_{C^{1},p}\lesssim f(q)|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\kappa},p}\lesssim(p-1)^{\frac{1}{2}}L\lambda_{q+1}^{\frac{\gamma}{8}},$
$|\!|\!|z_{q}|\!|\!|_{C^{\frac{1}{2}-\delta}_{t}C^{0},p}$ $\lesssim|\!|\!|\tilde{z}_{q}|\!|\!|_{C^{\frac{1}{2}-\delta}_{t}C^{1}_{x},2p}|\!|\!|\tilde{z}_{q}|\!|\!|_{C^{\frac{1}{2}-\delta}_{t}C^{0},2p}\lesssim\lambda_{q+1}^{\frac{\gamma}{8}}|\!|\!|z|\!|\!|_{C^{\frac{1}{2}-\delta}_{t}H^{\frac{3}{2}+\kappa},2p}^{2}\lesssim(2p-1)L^{2}\lambda_{q+1}^{\frac{\gamma}{8}}.$
The above estimates (3.7) will be frequently used in the sequel. Under the above assumptions, our main iteration reads as follows, and the proof of this result is presented in Section 4 below.
Proposition 3.2.
Suppose that $\mathrm{Tr}((-\Delta)^{\frac{3}{2}+\sigma}GG^{*})<\infty$ for some $\sigma>0$ and let $r>1$ be fixed. Given a smooth function $e:\mathbb{R}\to(0,\infty)$ such that $\bar{e}\geqslant e(t)\geqslant\underline{e}\geqslant 6{\cdot 48}{\cdot(2\pi)^{3}}c_{R}^{-1}L^{2}$ with $\|e^{\prime}\|_{C^{0}}\leqslant\tilde{e}$ for some constants $\overline{e},\underline{e},\tilde{e}>0$, where $L$ is from Proposition 3.2 and $c_{R}>0$ is a small universal constant (see Section 4.3). There exists a choice of parameters $a,b,\beta$ and $\gamma$ such that the following holds true: Let $(v_{q},\mathring{R}_{q})$ for some $q\in\mathbb{N}_{0}$ be an $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted solution to (3.4) satisfying
$\|v_{q}\|_{C_{t,x}^{0}}$ $\leqslant\lambda_{q+2}^{\gamma},$ (3.8a)
$|\!|\!|v_{q}|\!|\!|_{C^{0},2r}$ $\leqslant\lambda_{q}^{\beta},$ (3.8b)
$\|v_{q}\|_{C_{t,x}^{1}}$ $\leqslant\lambda^{\frac{3}{2}}_{q}\delta_{q}^{\frac{1}{2}},$ (3.8c)
$\|\mathring{R}_{q}\|_{C^{0}_{t,x}}$ $\leqslant\lambda^{\gamma}_{q+3},$ (3.8d)
$|\!|\!|\mathring{R}_{q}|\!|\!|_{C^{0},r}$ $\leqslant\delta_{q+1},$ (3.8e)
$|\!|\!|\mathring{R}_{q}|\!|\!|_{L^{1},1}$ $\leqslant\frac{1}{48}c_{R}\delta_{q+2}\underline{e},$ (3.8f)
Moreover, for any $t\in{\mathbb{R}}$
$\frac{3}{4}\delta_{q+1}e(t)\leqslant e(t)-\mathbf{E}\|(v_{q}+z_{q})(t)\|_{L^{2}}^{2}\leqslant\frac{5}{4}\delta_{q+1}e(t),\quad$ $q\geqslant 1,$ (3.9)
$\mathbf{E}\|v_{0}(t)+z_{0}(t)\|^{2}_{L^{2}}\leqslant\frac{1}{2}e(t),\quad$ $q=0.$
Then there exists an $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted process $(v_{q+1},\mathring{R}_{q+1})$ which solves (3.4), obeys (3.8a)-(3.8f) and (3.9) at the level $q+1$ and satisfies
$|\!|\!|v_{q+1}-v_{q}|\!|\!|_{C^{0},2r}\leqslant\bar{M}\delta_{q+1}^{\frac{1}{2}},$ (3.10)
where $\bar{M}$ is a universal constant which will be fixed throughout the iteration.
We intend to start the iteration from $v_{0}\equiv 0$ on ${\mathbb{R}}$. In this case, we have $\mathring{R}_{0}=z_{0}\mathring{\otimes}z_{0}-\mathcal{R}z_{0}$. Recalling the definition of $z_{0}$ in (3.5), we obtain
$\|\mathring{R}_{0}\|_{C^{0}_{t,x}}\leqslant\|z_{0}\|^{2}_{C^{0}_{t,x}}+\|z_{0}\|_{C^{0}_{t,x}}\leqslant\lambda_{3}^{\frac{1}{2}\gamma}+\lambda_{3}^{\frac{1}{4}\gamma}\leqslant\lambda_{3}^{\gamma},$
where we choose $a$ sufficiently large to absorb the constant in the last inequality. Taking expectation and using (3.7), we obtain
$|\!|\!|\mathring{R}_{0}|\!|\!|_{C^{0},r}\leqslant|\!|\!|z_{0}|\!|\!|^{2}_{C^{0},2r}+|\!|\!|z_{0}|\!|\!|_{C^{0},r}\leqslant|\!|\!|z|\!|\!|^{2}_{H^{\frac{3}{2}+\sigma},2r}+|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},r}\leqslant 2rL^{2}+rL\leqslant\delta_{1},$
$|\!|\!|\mathring{R}_{0}|\!|\!|_{L^{1},1}\leqslant(2\pi)^{3}|\!|\!|\mathring{R}_{0}|\!|\!|_{C^{0},1}\leqslant 3{\cdot}(2\pi)^{3}L^{2}\leqslant\frac{1}{96}c_{R}\underline{e}\leqslant\frac{1}{48}c_{R}\delta_{2}e(t),$
where we used $6{\cdot}48{\cdot}(2\pi)^{3}c_{R}^{-1}L^{2}\leqslant\underline{e}$, $\delta_{1}=3rL^{2}$ and $\delta_{2}=\frac{1}{2}$. Hence, (3.8d), (3.8e) and (3.8f) are satisfied on the level $q=0$. Similarly, (3.9) also holds at the level $q=0$, since we have
$\mathbf{E}\|z_{0}(t)\|^{2}_{L^{2}}\leqslant(2\pi)^{3}\mathbf{E}\|z_{0}(t)\|^{2}_{C^{0}}\leqslant(2\pi)^{3}\mathbf{E}\|z(t)\|^{2}_{H^{\frac{3}{2}+\sigma}}\leqslant(2\pi)^{3}L^{2}\leqslant\frac{1}{2}e(t).$
We will see in the following context, as long as the $q=0$ step in (3.9) holds ture, $\zeta_{0}(t)$ in (4.11) below is well-defined, thus the iteration can proceed.
3.2. Proof of main result
Since the first iteration is established, we deduce the following result by using Proposition 3.2.
Theorem 3.3.
Suppose that $\mathrm{Tr}((-\Delta)^{\frac{3}{2}+\sigma}GG^{*})<\infty$ for some $\sigma>0$. Let $r>1$ be fixed and given a smooth function $e:\mathbb{R}\to(0,\infty)$ such that $\bar{e}\geqslant e(t)\geqslant\underline{e}\geqslant 6{\cdot 48}{\cdot(2\pi)^{3}}c_{R}^{-1}L^{2}$ with $\|e^{\prime}\|_{C^{0}}\leqslant\tilde{e}$ for some constants $\overline{e},\underline{e},\tilde{e}>0$. There exists an $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted process $u$ which belongs to $C({\mathbb{R}};C^{\vartheta})$ $\mathbf{P}$-a.s. for some $\vartheta>0$ and is an analytically weak solution to (1.1) in the sense of Definition 1.1. Moreover, the solution satisfies
$|\!|\!|u|\!|\!|_{C^{\vartheta},2r}<\infty,$ (3.11)
and for all $t\in{\mathbb{R}}$
$\mathbf{E}\|u(t)\|_{L^{2}}^{2}=e(t).$ (3.12)
There are infinitely many such solutions by choosing different energies $e$.
Proof.
We may thus start the iteration Proposition 3.2 with the pair $(v_{0},\mathring{R}_{0})$ and obtain a sequence of solutions $(v_{q},\mathring{R}_{q})$. By (3.8c), (3.10) and interpolation we deduce for any $\vartheta\in\left(0,{\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot}b^{5}},\frac{1}{21{\cdot}b^{4}}\}}\right)\in(0,\frac{2}{3}\beta)$, the following series is summable
$\sum_{q\geqslant 0}|\!|\!|v_{q+1}-v_{q}|\!|\!|_{C^{\vartheta},2r}$ $\leqslant\sum_{q\geqslant 0}|\!|\!|v_{q+1}-v_{q}|\!|\!|^{1-\vartheta}_{C^{0},2r}|\!|\!|v_{q+1}-v_{q}|\!|\!|^{\vartheta}_{C^{1},2r}$
$\lesssim\sum_{q\geqslant 0}\delta_{q+1}^{\frac{1-\vartheta}{2}}\lambda_{q+1}^{\frac{3}{2}\vartheta}\delta_{q+1}^{\frac{1}{2}\vartheta}\leqslant\sqrt{3r}La^{\frac{3}{2}b\vartheta}+\lambda_{2}^{\beta}\sum_{q\geqslant 1}\lambda_{q+1}^{\frac{3}{2}\vartheta-\beta}<\infty.$
Thus, we may define a limiting function $v=\lim_{q\to\infty}v_{q}$ which lies in $L^{2r}(\Omega;C({\mathbb{R}},C^{\vartheta}))$. Since $v_{q}$ is $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted for every $q\in\mathbb{N}_{0}$, the limit $v$ is $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted as well. Combining (3.7) with (4.33) in the proof of Proposition 3.2 and using interpolation again, we deduce for the same $\vartheta$ as above and any $p\geqslant 1$
$\sum_{q\geqslant 0}|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{\vartheta},p}$ $\leqslant\sum_{q\geqslant 0}|\!|\!|z_{q+1}-z_{q}|\!|\!|^{1-\vartheta}_{C^{0},p}|\!|\!|z_{q+1}-z_{q}|\!|\!|^{\vartheta}_{C^{1},p}$
$\lesssim p\sum_{q\geqslant 0}(\lambda_{q+1}^{-\frac{\gamma}{8}\sigma}+\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{8}})^{1-\vartheta}\lambda_{q+2}^{\frac{\gamma}{8}\vartheta}$
$\lesssim p\sum_{q\geqslant 1}\left(\lambda_{q+1}^{\frac{\gamma}{8}(b\vartheta+\sigma\vartheta-\sigma)}+\lambda_{q+1}^{\gamma(\frac{9}{8}b\vartheta+\frac{1}{8}-b)}\right)<\infty,$
since $\vartheta<{\mathord{{\rm min}}}{\{\frac{\sigma}{b+\sigma},\frac{8b-1}{9b}\}}$. Then we obtain $\lim_{q\to\infty}z_{q}=z$ in $L^{p}(\Omega;C({\mathbb{R}},C^{\vartheta}))$ for any $p\geqslant 1$. Furthermore, it follows from (3.8e) that $\lim_{q\to\infty}\mathring{R}_{q}=0$ in $L^{1}(\Omega;C({\mathbb{R}},C^{0}))$. Thus $v$ is an analytically weak solution to (3.2). Hence letting $u=v+z$ we obtain an $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$-adapted analytically weak solution to (1.1). Moreover, the estimate (3.11) for $u$ holds. Finally, (3.12) follows from (3.9), and choosing different energy functions, we can obtain infinitely many such solutions. This concludes the proof of Theorem 3.3. ∎
4. Proof of Proposition 3.2
The proof proceeds in several main steps which are similar as in many convex integration schemes. First of all, we start the construction by fixing the parameters in Section 4.1 and proceed with a mollification step in Section 4.2. Section 4.3 is devoted to introduce the new iteration $v_{q+1}$. This is the main part of the construction which differs in general convex integration schemes since we aim to construct global Hölder continuous flows. Here, we combine the construction methods of [BV19] and [HZZ22b] to obtain a new perturbation $w_{q+1}$. More precisely, we construct new amplitude functions $a_{(\xi)}$ similarly to [HZZ22b], meanwhile, we incorporate the solutions for transport equations into Beltrami waves to obtain an acceptable transport error similar to [BV19]. As pathwise positive lower bounds of the solutions to the transport equations are required, we use cutoff method to control the growth of noise and add pathwise estimtes of the velocity field $v_{q}$ in inductive iteration. Section 4.4 contains the inductive estimates of $v_{q+1}$, especially the pathwise and moment estimates. In Section 4.5, we show how to construct the new stress $\mathring{R}_{q+1}$. We establish the inductive pathwise and moment estimates on $\mathring{R}_{q+1}$ in Section 4.6. Finally, in Section 4.7 we show how the energy is controlled.
4.1. Choice of parameters
In the sequel, additional parameters will be indispensable and their value has to be carefully chosen in order to respect all the compatibility conditions appearing in the estimations below. First, for a fixed integer $b\geqslant 7$, we let $\beta,\gamma\in(0,1)$ be small parameters satisfying
$4b^{3}\gamma+4b^{3}\beta<1,\qquad 3b\beta<\gamma.$
In the sequel, we also use the following bound
$2b^{2}\beta+\beta<\frac{\gamma\sigma}{8},$
which can be obtained by choosing $\gamma=\frac{2}{9b^{3}}$ and choosing $\beta$ small such that $\beta<{\mathord{{\rm min}}}{\{\frac{\sigma}{78{\cdot}b^{5}},\frac{2}{27{\cdot}b^{4}}\}}$. The last free parameter $a\in 2^{36b^{2}\mathbb{N}}$ is chosen sufficiently large such that $a^{\frac{b\gamma}{8}}\in\mathbb{N}$ and $2<a^{(\frac{b}{4}-1)b^{2}\gamma}$. The choice of $a$ ensures $f(q)\in\mathbb{N}$ and (3.6) holds. The above requirements for parameters will be used in the following sections. In the sequel, we increase $a$ in order to absorb various implicit and universal constants.
4.2. Mollification
We replace $v_{q}$ by a mollified velocity field $v_{\ell}$, we choose a small parameter $\ell\in(0,1)$ as $\ell=\lambda_{q}^{-2}$. Let $\{\phi_{\varepsilon}\}_{\varepsilon>0}$ be a family of mollifiers on $\mathbb{R}^{3}$, and $\{\varphi_{\varepsilon}\}_{\varepsilon>0}$ be a family of mollifiers with support in $(0,1)$. The one-sided mollifiers here is used in order to preserve adaptedness. We define a mollification of $v_{q}$ and $\mathring{R}_{q}$ in space and time by
$v_{\ell}=(v_{q}*_{x}\phi_{\ell})*_{t}\varphi_{\ell},\qquad\mathring{R}_{\ell}=(\mathring{R}_{q}*_{x}\phi_{\ell})*_{t}\varphi_{\ell},\qquad z_{\ell}=({z_{q}}*_{x}\phi_{\ell})*_{t}\varphi_{\ell},$
where $\phi_{\ell}=\frac{1}{\ell^{3}}\phi(\frac{\cdot}{\ell})$ and $\varphi_{\ell}=\frac{1}{\ell}\varphi(\frac{\cdot}{\ell})$. By definition it is easy to know that $z_{\ell}$ is adapted. Then using (3.4) we obtain that $(v_{\ell},\mathring{R}_{\ell})$ satisfies
$\partial_{t}v_{\ell}-z_{\ell}+\mathord{{\rm div}}((v_{\ell}+z_{\ell})\otimes(v_{\ell}+z_{\ell}))+\nabla p_{\ell}$ $=\mathord{{\rm div}}(\mathring{R}_{\ell}+R_{\textrm{com}}),$ (4.1)
$\mathord{{\rm div}}v_{\ell}$ $=0,$
where
$R_{\textrm{com}}=(v_{\ell}+z_{\ell})\mathring{\otimes}(v_{\ell}+z_{\ell})-(((v_{q}+z_{q})\mathring{\otimes}(v_{q}+z_{q}))*_{x}\phi_{\ell})*_{t}\varphi_{\ell},$ (4.2)
$p_{\ell}=(p_{q}*_{x}\phi_{\ell})*_{t}\varphi_{\ell}-\frac{1}{3}\big{(}|v_{\ell}+z_{\ell}|^{2}-(|v_{q}+z_{q}|^{2}*_{x}\phi_{\ell})*_{t}\varphi_{\ell}).$
Using the definition of $z_{q}$ (3.5) and the pathwise inductive assumption (3.8a), we can obtain the following pathwise estimate for $R_{\textrm{com}}$
$\|R_{\textrm{com}}\|_{C^{0}_{t,x}}\leqslant 2\|v_{q}+z_{q}\|^{2}_{C^{0}_{[t-1,t+1],x}}\leqslant 4(\lambda_{q+2}^{2\gamma}+\lambda_{q+3}^{\frac{\gamma}{2}})\leqslant\frac{1}{10}\lambda_{q+4}^{\gamma},$ (4.3)
where we choose $a$ sufficiently large to absorb the constant in the last inequality. Using a mollification estimate and inductive assumption (3.8a) and (3.8c), for any $t\in\mathbb{R}$ and $\delta\in(0,\frac{1}{12})$, we have
$\|R_{\textrm{com}}(t)\|_{L^{\infty}}$
$\lesssim\left(\ell\|v_{q}\|_{C_{[t-1,t],x}^{1}}+\ell\|z_{q}\|_{C^{0}_{[t-1,t]}C_{x}^{1}}+\ell^{\frac{1}{2}-\delta}\|z_{q}\|_{C^{\frac{1}{2}-\delta}_{[t-1,t]}C_{x}^{0}}\right)\left(\|z_{q}\|_{C^{0}_{[t-1,t],x}}+\|v_{q}\|_{C^{0}_{[t-1,t],x}}\right).$
Taking expectation and using (3.7), (3.8a) and (3.8b), we obtain
$|\!|\!|\mathring{R}_{\textrm{com}}|\!|\!|_{C^{0},r}$ $\lesssim\left(\ell\lambda_{q}^{\frac{3}{2}}+\ell|\!|\!|z_{q}|\!|\!|_{C^{1},2r}+\ell^{\frac{1}{2}-\delta}|\!|\!|z_{q}|\!|\!|_{C^{\frac{1}{2}-\delta}_{t}C^{0},2r}\right)\left(|\!|\!|z_{q}|\!|\!|_{C^{0},2r}+|\!|\!|v_{q}|\!|\!|_{C^{0},2r}\right)$ (4.4)
$\lesssim\left(\ell\lambda_{q}^{\frac{3}{2}}+\ell\lambda_{q+1}^{\frac{\gamma}{8}}2rL+\ell^{\frac{1}{2}-\delta}\lambda_{q+1}^{\frac{\gamma}{8}}2pL^{2}\right)\left(\lambda_{q}^{\beta}+2rL\right)$
$\lesssim\lambda_{q}^{-\frac{1}{2}+\beta}+\lambda_{q}^{-\frac{3}{4}+\beta}\lambda_{q+1}^{\frac{\gamma}{8}}$
$\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\leqslant\frac{1}{6}\delta_{q+2},$
where we used $b\gamma+4b^{3}\beta<1$ to have $\lambda_{q}^{-\frac{1}{2}+\beta}<\lambda_{q}^{-2b^{3}\beta}$ and $\lambda_{q}^{-\frac{3}{4}+\beta+\frac{1}{8}b\gamma}<\lambda_{q}^{-2b^{3}\beta}$ and $a$ was chosen sufficiently large to absorb the constant. By (4.4), we also obtain
$\qquad|\!|\!|\mathring{R}_{\textrm{com}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$ (4.5)
Similarly, by inductive assumption (3.8c) and the same requirements for parameters as above, we also obtain for any $t\in\mathbb{R}$
$\|(v_{q}-v_{\ell})(t)\|_{L^{\infty}}\lesssim\ell\|v_{q}\|_{C_{[t-1,t],x}^{1}}\leqslant\ell\lambda_{q}^{\frac{3}{2}}\delta_{q}^{\frac{1}{2}}\leqslant\frac{1}{4}\delta_{q+1}^{\frac{1}{2}}.$ (4.6)
4.3. Perturbation of the velocity
As usual in convex integration schemes, in order to guarantee smoothness throughout the construction, the new velocity field $v_{q+1}$ will be constructed as a perturbation of $v_{\ell}$:
$v_{q+1}=:v_{\ell}+w_{q+1},$
where $w_{q+1}:=w_{q+1}^{(p)}+w_{q+1}^{(c)}$ is constructed following [BV19]. We proceed construction with introducing transport equations and time cutoffs.
4.3.1. Flow maps and cutoffs
In order to have an acceptable transport error in the estimates of $\mathring{R}_{q+1}$, the perturbation $w_{q+1}$ needs to be transported by the flow of the vector field $\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla$. Similar to [BV19], a natural way to achieve this, is to replace the linear phase $\xi\cdot x$ in the definition of the Beltrami wave $W_{\xi,\lambda}$ by the nonlinear phase $\xi\cdot\Phi(x,t)$, where $\Phi$ is transported by the aforementioned vector field. A slight difference from [BV19] is that we extend the definition of $\Phi$ to the whole line w.r.t time variable. For any integer $k\in\mathbb{Z}$, we subdivide $[k,k+1]$ into time intervals of size $\ell$ and solve transport equations on these intervals. For $j\in\{0,1,...,\lceil\ell^{-1}\rceil\}$ we define the adapted map $\Phi_{k,j}:\Omega\times{{\mathbb{R}}^{3}}\times[k+(j-1)\ell,k+(j+1)\ell]\to{{\mathbb{R}}^{3}}$ as the ${\mathbb{T}}^{3}$ periodic solution of
$(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)\Phi_{k,j}$ $=0,$ (4.7)
$\Phi_{k,j}(k+(j-1)\ell,x)$ $=x.$
We put further details on the $C_{x}^{n}$-estimates of $\Phi_{k,j}$ in Appendix C. Specifically, the following two estimates which follows from (C.5) and (C.6)
$\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)-\textrm{Id}\|_{C_{x}^{0}}\leqslant\lambda^{-\frac{1}{2}}_{q}\ll 1,$ (4.8a)
$\frac{1}{2}\leqslant$ $\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)\|_{C_{x}^{0}}\leqslant 2.$ (4.8b)
As indicated by the proof in Appendix C, the validity of (4.8a) requires that both $v_{q}$ and $z_{q}$ have pointwise bounds, and this is why we need to cut off $\tilde{z}_{q}$ and introduce (3.8c) in the inductive iteration, instead of solely considering the moment bounds in the inductive iteration as in [HZZ22b]. The above estimates play a crucial role in the cancellation of the estimates for the new Reynolds stress $\mathring{R}_{q+1}$. In particular, (4.8b) ensures the validity of stationary phase Lemma 4.2 below in Section 4.6, which requires the positive lower bounds of $\nabla\Phi_{k,j}$. We could easily extend $\Phi_{k,j}$ from $\Omega\times\mathbb{R}^{3}\times[k+(j-1)\ell,k+(j+1)\ell]$ to $\Omega\times\mathbb{R}^{3}\times[k,k+1]$ in an adapted way, which is still denoted as $\Phi_{k,j}$. The extension is not unique but in the following we only use the value of $\Phi_{k,j}$ on $[k+(j-1)\ell,k+(j+1)\ell]$. We also let $\eta$ be a non-negative function supported in $(-1,1)$, which is 1 on $(-\frac{1}{4},\frac{1}{4})$ and such that the square of the shifted bump functions
$\eta_{j}(t)=\eta(\ell^{-1}t-j)$
form a partition of unity
$\sum_{j}\eta^{2}_{j}(t)=1,$
[…]
$\eta_{k,j}(t)=\eta(\ell^{-1}(t-k)-j)$
form a partition of unity
$\sum_{j}\eta^{2}_{k,j}(t)=1,$ (4.9)
for all $t\in[k,k+1]$.
4.3.2. Construction of $w_{q+1}$
Let us now proceed with the construction of the perturbation $w_{q+1}$. To this end, the building blocks of the perturbation are the Beltrami waves presented in [BV19, section 5.4], which we recall in Appendix A. Since we use moment bounds of $\mathring{R}_{q}$ to the iterative estimates in Proposition 3.2, we have to adjust the amplitude functions $a_{(\xi)}$ such that we can apply Lemma A.2 in our setting. More precisely, we first define $\rho$ as follows
$\rho(t,x):=\sqrt{\ell^{2}+|\mathring{R}_{\ell}(t,x)|^{2}}+c_{*}\zeta_{\ell}(t),$ (4.10)
$\zeta_{q}(t):=\frac{1}{3{\cdot}(2\pi)^{3}}\Big{[}e(t)(1-\delta_{q+2})-\mathbf{E}\|v_{q}(t)+{z_{q}}(t)\|_{L^{2}}^{2}\Big{]},$ (4.11)
and
$\zeta_{\ell}:=\zeta_{q}*_{t}\varphi_{\ell},$
where $c_{*}>0$ is a constant introduced in Lemma A.2 and we take the constant $c_{R}$ in (3.8f) satisfying $c_{R}<c_{*}$. Now, we define the amplitude functions
$a_{(\xi)}(t,x)=a_{q+1,j,\xi}(t,x)=c_{*}^{-\frac{1}{2}}\rho^{\frac{1}{2}}\eta_{k,j}(t)\gamma_{\xi}^{(j)}\left(\textrm{Id}-\frac{c_{*}\mathring{R}_{\ell}}{\rho}\right),\ t\in[k,k+1],$ (4.12)
where $\gamma_{\xi}^{(j)}$ is introduced in Appendix A. By the definition of $\rho$ we know
$\left\|\textrm{Id}-\left(\textrm{Id}-\frac{c_{*}\mathring{R}_{\ell}(t,x)}{\rho}\right)\right\|_{C^{0}}\leqslant c_{*}.$
[…]
$c_{*}^{-1}\rho\textrm{Id}-\mathring{R}_{\ell}=\frac{1}{2}\sum_{j}\sum_{\xi\in\Lambda_{j}}a^{2}_{(\xi)}(\textrm{Id}-\xi\otimes\xi)$ (4.13)
[…]
$\|a_{(\xi)}\|_{C^{N}_{t,x}}\lesssim\ell^{-2N-\frac{1}{2}}\lambda_{q+3}^{(N+1)\gamma},$ (4.14)
and
$\|a_{(\xi)}\|_{C^{0}_{t,x}}\lesssim\lambda_{q+3}^{\gamma},\qquad\|a_{(\xi)}\|_{C^{1}_{t,x}}\lesssim\ell^{-2}\lambda_{q+3}^{2\gamma}.$ (4.15)
With these preparations in hand, we define the principal part $w_{q+1}^{(p)}$ of the perturbation $w_{q+1}$ as in [BV19], more precisely, for $t\in[k,k+1]$, we define
$w_{(\xi)}(t,x):=a_{q+1,j,\xi}(t,x)B_{\xi}e^{i\lambda_{q+1}\xi\cdot\Phi_{k,j}(t,x)},$
and
$w^{(p)}_{q+1}(t,x):$ $=\sum_{j}\sum_{\xi\in\Lambda_{j}}a_{q+1,j,\xi}(t,x)B_{\xi}e^{i\lambda_{q+1}\xi\cdot\Phi_{k,j}(t,x)}$ (4.16)
$=\sum_{j}\sum_{\xi\in\Lambda_{j}}c_{*}^{-\frac{1}{2}}\rho^{\frac{1}{2}}\eta_{k,j}(t)\gamma_{\xi}^{(j)}\left(\textrm{Id}-\frac{c_{*}\mathring{R}_{\ell}}{\rho}\right)B_{\xi}e^{i\lambda_{q+1}\xi\cdot\Phi_{k,j}(t,x)},$
where $B_{\xi}$ is suitable vector defined in Appendix A. Since the coefficients $a_{(\xi)}$ and $\Phi_{k,j}(t,x)$ are $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted we deduce that $w_{q+1}^{(p)}$ is $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted. Next, we define the incompressibility correction $w_{q+1}^{(c)}$. We aim to add a corrector to $w_{q+1}^{(p)}$ such that the resulting function is perfect $\mathrm{curl}$, making it thus divergence free. To this end, it is useful to introduce the following scalar $phase$ function
$\phi_{(\xi)}(t,x):=e^{i\lambda_{q+1}\xi\cdot(\Phi_{k,j}(t,x)-x)},\quad t\in[k,k+1].$ (4.17)
Using the same notation as in Appendix A, we define
$W_{(\xi)}=B_{\xi}e^{i\lambda_{q+1}\xi\cdot x}.$
Since from Appendix A $\mathrm{curl}$ $W_{(\xi)}=\lambda_{q+1}W_{(\xi)}$, and $\phi_{(\xi)}$ and $a_{(\xi)}$ are scalar functions, we have
$a_{(\xi)}\phi_{(\xi)}W_{(\xi)}=\frac{1}{\lambda_{q+1}}\textrm{curl}\left(a_{(\xi)}\phi_{(\xi)}W_{(\xi)}\right)-\frac{1}{\lambda_{q+1}}\nabla\left(a_{(\xi)}\phi_{(\xi)}\right)\times W_{(\xi)}.$
We therefore define
$w_{(\xi)}^{(c)}(t,x):$ $=\frac{1}{\lambda_{q+1}}\nabla(a_{(\xi)}\phi_{(\xi)})\times B_{\xi}e^{i\lambda_{q+1}\xi\cdot x}$ (4.18)
$=\left(\frac{\nabla a_{(\xi)}}{\lambda_{q+1}}+ia_{(\xi)}(\nabla\Phi_{k,j}(t,x)-\textrm{Id})\xi\right)\times W_{(\xi)}(\Phi_{k,j}(t,x)),\ t\in[k.k+1].$
The incompressibility correction $w^{(c)}_{q+1}$ is defined as
$w^{(c)}_{q+1}(t,x)=\sum_{j}\sum_{\xi\in\Lambda_{j}}w_{(\xi)}^{(c)}(t,x)$
for all $t\in[k,k+1]$, $k\in\mathbb{Z}$. Since the coefficients $a_{(\xi)}$ and $\Phi_{k,j}(t,x)$ are $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted we deduce that $w_{q+1}^{(c)}$ is $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted, and the new perturbation is defined as
$w_{q+1}=w^{(p)}_{q+1}+w^{(c)}_{q+1}=\frac{1}{\lambda_{q+1}}\sum_{j}\sum_{\xi\in\Lambda_{j}}\textrm{curl}\left(a_{(\xi)}\phi_{(\xi)}W_{(\xi)}\right),$ (4.19)
and so clearly $w_{q+1}$ is mean zero, divergence free and $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted. We define the velocity field at level $q+1$ as
$v_{q+1}:=v_{\ell}+w_{q+1}.$
Thus, it is also $(\mathcal{F}_{t})_{t\geqslant 0}$-adapted.
4.4. Inductive estimates for $v_{q+1}$
In this section, we verify the inductive estimates (3.8a), (3.8b), (3.8c) and (3.10). By the definition of $w^{(p)}_{q+1}$ (4.16), it follows from (3.8d), (4.10) and (A.5) that
$\|w^{(p)}_{q+1}\|_{C^{0}_{t,x}}\lesssim 2|\Lambda_{j}|c_{*}^{-\frac{1}{2}}M(\|\mathring{R}_{q}\|^{\frac{1}{2}}_{C^{0}_{[t-1,t+1],x}}+\ell^{\frac{1}{2}})+2|\Lambda_{j}|M\bar{e}\delta_{q+1}^{\frac{1}{2}}\leqslant\lambda_{q+3}^{\frac{3}{4}\gamma},$ (4.20)
where $M$ is the universal constant in (A.5). Taking expectation on the both side of (4.20), we then use (3.8e) to obtain
$|\!|\!|w^{(p)}_{q+1}|\!|\!|_{C^{0},2r}\lesssim 2|\Lambda_{j}|c_{*}^{-\frac{1}{2}}M(|\!|\!|\mathring{R}_{q}|\!|\!|_{C^{0},r}^{\frac{1}{2}}+\ell^{\frac{1}{2}})+2|\Lambda_{j}|M\bar{e}\delta_{q+1}^{\frac{1}{2}}\leqslant\frac{1}{4}\bar{M}\delta_{q+1}^{\frac{1}{2}},$ (4.21)
where $\bar{M}$ is a universal constant satisfying $9c_{*}^{-\frac{1}{2}}M|\Lambda_{j}|+8M\bar{e}|\Lambda_{j}|<\bar{M}$ and we choose $a$ sufficiently large to absorb the constant in the above two inequalities. Applying (4.8b), (4.15) and (C.8), we also obtain
$\|w^{(p)}_{q+1}\|_{C^{1}_{t,x}}$ $\lesssim\sup_{j}\sum_{\xi\in\Lambda_{j}}\|a_{(\xi)}\|_{C^{1}_{t,x}}+\lambda_{q+1}\|a_{(\xi)}\|_{C^{0}_{t,x}}\left(\|\partial_{t}\Phi_{k,j}\|_{C^{0}}+\|\nabla\Phi_{k,j}\|_{C^{0}}\right)$ (4.22)
$\lesssim\lambda_{q}^{4}\lambda_{q+3}^{2\gamma}+\lambda_{q+3}^{2\gamma}\lambda_{q+1}\leqslant\frac{1}{4}\lambda_{q+1}^{\frac{3}{2}}\delta_{q+1}^{\frac{1}{2}},$
where we used $b\geqslant 7$ and $12b^{3}\gamma+b\beta<$ 1 to have $\lambda_{q}^{2b^{3}\gamma+4}<\lambda_{q}^{\frac{3}{2}b-b\beta}$ and $\lambda_{q}^{2b^{3}\gamma}<\lambda_{q}^{\frac{1}{2}b-b\beta{\tiny}}$ in the last inequality, and $a$ was chosen sufficiently large to absorb the constant. We proceed to estimate $w^{(c)}_{q+1}$, using a standard mollification estimate, (4.8a) and (4.15) implies
$\|w^{(c)}_{q+1}\|_{C^{0}_{t,x}}$ $\lesssim\sup_{j}\sum_{\xi\in\Lambda_{j}}\left(\frac{\|\nabla a_{(\xi)}\|_{C^{0}_{t,x}}}{\lambda_{q+1}}+\|a_{(\xi)}\|_{C^{0}_{t,x}}\cdot\|\nabla\Phi_{k,j}-\textrm{Id}\|_{C^{0}}\right)$ (4.23)
$\lesssim\frac{\lambda_{q}^{4}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}}+\lambda_{q+3}^{\gamma}\lambda_{q}^{-\frac{1}{2}}\leqslant\frac{1}{4}\delta_{q+1}^{\frac{1}{2}},$
where we used $b\geqslant 7$ and $12b^{3}\gamma+2b\beta<$ 1 to have $\lambda_{q}^{b^{3}\gamma}<\lambda_{q}^{\frac{1}{2}-b\beta}$ and $\lambda_{q}^{4+2b^{3}\gamma}<\lambda_{q}^{b-b\beta}$ in the last inequality, and $a$ was chosen sufficiently large to absorb the constant. Taking expectation we obtain
$|\!|\!|w_{q+1}^{(c)}|\!|\!|_{C^{0},2r}\leqslant\frac{1}{4}\delta_{q+1}^{\frac{1}{2}}.$ (4.24)
Using (4.8a), (4.8b), (4.14), (4.15), (4.23) and estimates (C.7)–(C.9) for $\Phi_{k,j}$, we obtain
$\|w^{(c)}_{q+1}\|_{C_{t,x}^{1}}\lesssim$ $\lambda_{q+1}\|w^{(c)}_{q+1}\|_{C^{0}_{t,x}}\sup_{j}(\|\partial_{t}\Phi_{k,j}\|_{C^{0}}+\|\nabla\Phi_{k,j}\|_{C^{0}})$ (4.25)
$+\sup_{j}\sum_{\xi\in\Lambda_{j}}\bigg{[}\frac{\|\nabla a_{(\xi)}\|_{C^{1}_{t,x}}}{\lambda_{q+1}}+\|a_{(\xi)}\|_{C^{1}_{t,x}}\|\nabla\Phi_{k,j}-\textrm{Id}\|_{C^{0}}$
$\qquad\qquad\quad+\|a_{(\xi)}\|_{C^{0}_{t,x}}(\|\nabla^{2}\Phi_{k,j}\|_{C^{0}}+\|\partial_{t}\nabla\Phi_{k,j}\|_{C^{0}})\bigg{]}$
$\lesssim$ $\lambda_{q+1}\lambda_{q+3}^{\gamma}\delta_{q+1}^{\frac{1}{2}}+\frac{\lambda_{q}^{9}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}}+\lambda_{q}^{4}\lambda_{q+3}^{2\gamma}\lambda_{q}^{-\frac{1}{2}}+\lambda_{q+3}^{\gamma}(\ell^{-\frac{3}{4}}+\lambda_{q}^{\frac{3}{2}}\lambda_{q+3}^{\gamma})$
$\leqslant$ $\lambda_{q+1}\lambda_{q+3}^{\gamma}\delta_{q+1}^{\frac{1}{2}}+\lambda_{q}^{4}\lambda_{q+3}^{2\gamma}+\frac{\lambda_{q}^{9}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}}\leqslant\frac{1}{4}\lambda_{q+1}^{\frac{3}{2}}\delta_{q+1}^{\frac{1}{2}},$
where we used $b\geqslant 7$ and $12b\gamma^{3}+b\beta<$ 1 to have $\lambda_{q}^{b^{3}\gamma}<\lambda_{q}^{\frac{b}{2}}$, $\lambda_{q}^{4+2b^{3}\gamma}<\lambda_{q}^{\frac{3}{2}b-b\beta}$ and $\lambda_{q}^{9+3b^{3}\gamma}<\lambda_{q}^{\frac{5}{2}b-b\beta}$ in the last inequality, and $a$ was chosen sufficiently large to absorb the constant.
By (4.21) and (4.24), we have
$|\!|\!|w_{q+1}|\!|\!|_{C^{0},2r}\leqslant\frac{3}{4}\bar{M}\delta_{q+1}^{\frac{1}{2}},$
[…]
$|\!|\!|v_{\ell}-v_{q}|\!|\!|_{C^{0},2r}\leqslant\frac{1}{4}\delta_{q+1}^{\frac{1}{2}}.$
Combining the above two estimates, we obtain
$|\!|\!|v_{q+1}-v_{q}|\!|\!|_{C^{0},2r}\leqslant|\!|\!|v_{\ell}-v_{q}|\!|\!|_{C^{0},2r}+|\!|\!|w_{q+1}|\!|\!|_{C^{0},2r}\leqslant\bar{M}\delta_{q+1}^{\frac{1}{2}}.$
Hence, (3.10) holds at the level $q+1$. Moreover, it follows from (3.8a), (4.20) and (4.23) that
$\|v_{q+1}\|_{C^{0}_{t,x}}\leqslant\|v_{\ell}\|_{C^{0}_{t,x}}+\|w_{q+1}^{(p)}\|_{C^{0}_{t,x}}+\|w_{q+1}^{(c)}\|_{C^{0}_{t,x}}\leqslant\lambda_{q+2}^{\gamma}+\lambda_{q+3}^{\frac{3}{4}\gamma}+\delta_{q+1}^{\frac{1}{2}}\leqslant\lambda_{q+3}^{\gamma}.$ (4.26)
By (3.8c), (4.22) and (4.25) we obtain
$\|v_{q+1}\|_{C^{1}_{t,x}}\leqslant\|v_{\ell}\|_{C^{1}_{t,x}}+\|w_{q+1}^{(p)}\|_{C^{1}_{t,x}}+\|w_{q+1}^{(c)}\|_{C^{1}_{t,x}}\leqslant\lambda^{\frac{3}{2}}_{q}\delta^{\frac{1}{2}}_{q}+\frac{1}{2}\lambda^{\frac{3}{2}}_{q+1}\delta^{\frac{1}{2}}_{q+1}\leqslant\lambda^{\frac{3}{2}}_{q+1}\delta^{\frac{1}{2}}_{q+1}.$
By (3.8b), (4.21) and (4.24) we obtain
$|\!|\!|v_{q+1}|\!|\!|_{C^{0},2r}\leqslant|\!|\!|v_{q}|\!|\!|_{C^{0},2r}+|\!|\!|w^{(p)}_{q+1}|\!|\!|_{C^{0},2r}+|\!|\!|w^{(c)}_{q+1}|\!|\!|_{C^{0},2r}\leqslant\lambda_{q}^{\beta}+\bar{M}\delta_{q+1}^{\frac{1}{2}}\leqslant\lambda_{q+1}^{\beta}.$ (4.27)
Therefore, we have verified that (3.8a), (3.8b) and (3.8c) hold at the level $q+1$.
4.5. Definition of the Reynolds stress $\mathring{R}_{q+1}$
[…]
$\mathord{{\rm div}}\mathring{R}_{q+1}-\nabla p_{q+1}$ (4.28)
$=\underbrace{(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(p)}}_{\mathord{{\rm div}}(R_{\textrm{transport}})}+\underbrace{\mathord{{\rm div}}(w_{q+1}^{(p)}\otimes w_{q+1}^{(p)}+\mathring{R}_{\ell})}_{\mathord{{\rm div}}(R_{\textrm{oscillation}})+\nabla p_{\textrm{oscillation}}}$
$+\underbrace{w_{q+1}\cdot\nabla(v_{\ell}+z_{\ell})}_{\mathord{{\rm div}}(R_{\textrm{Nash}})}+\underbrace{(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(c)}+\mathord{{\rm div}}(w_{q+1}^{(c)}\otimes w_{q+1}+w_{q+1}^{(p)}\otimes w_{q+1}^{(c)})}_{\mathord{{\rm div}}(R_{\textrm{corrector}})+\nabla p_{\textrm{corrector}}}$
$+\underbrace{\mathord{{\rm div}}(v_{q+1}\otimes z_{q+1}-v_{q+1}\otimes z_{\ell}+z_{q+1}\otimes v_{q+1}-z_{\ell}\otimes v_{q+1}+z_{q+1}\otimes z_{q+1}-z_{\ell}\otimes z_{\ell})-z_{q+1}+z_{\ell}}_{\mathord{{\rm div}}(R_{\textrm{commutator1}})+\nabla p_{\textrm{commutator1}}}$
$+\mathord{{\rm div}}(R_{\textrm{commutator}})-\nabla p_{\ell}.$
[…]
$R_{\textrm{transport}}$ $:=\mathcal{R}\left((\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(p)}\right),$
$R_{\textrm{Nash}}$ $:=\mathcal{R}(w_{q+1}\cdot\nabla(v_{\ell}+z_{\ell})),$
$R_{\textrm{corrector}}$ $:=\mathcal{R}\left((\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(c)}\right)+\left(w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(c)}+w_{q+1}^{(p)}\mathring{\otimes}w_{q+1}^{(c)}+w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(p)}\right),$
$R_{\textrm{commutator1}}$ $:=v_{q+1}\mathring{\otimes}(z_{q+1}-z_{\ell})+(z_{q+1}-z_{\ell})\mathring{\otimes}v_{q+1}+z_{q+1}\mathring{\otimes}z_{q+1}-z_{\ell}\mathring{\otimes}z_{\ell}-\mathcal{R}(z_{q+1}-z_{\ell}),$
$\nabla p_{\textrm{corrector}}$ $:=\frac{1}{3}(2w_{q+1}^{(c)}\cdot w_{q+1}^{(p)}+|w_{q+1}^{(c)}|^{2}),$
$\nabla p_{\textrm{commutator1}}$ $:=\frac{1}{3}(v_{q+1}\cdot z_{q+1}-v_{q+1}\cdot z_{\ell}+z_{q+1}\cdot v_{q+1}-z_{\ell}\cdot v_{q+1}+z_{q+1}\cdot z_{q+1}-z_{\ell}\cdot z_{\ell}).$
In order to define the remaining oscillation error from the second line in (4.28), first note that for $j,j^{\prime}$ such that $|j-j^{\prime}|\geqslant 2$, we have $\eta_{k,j}(t)\eta_{k,j^{\prime}}(t)=0$ Second, we have $\Lambda_{j}\cap\Lambda_{j^{\prime}}=\emptyset$ for $|j-j^{\prime}|=1$, and by Lemma A.1 we have $\mathord{{\rm div}}(W_{\xi}\otimes W_{\xi^{\prime}}+W_{\xi^{\prime}}\otimes W_{\xi})=\nabla(W_{\xi}\cdot W_{\xi^{\prime}})$. Therefore, we may use (4.13) and Lemma A.2 to obtain
$\mathord{{\rm div}}(w_{q+1}^{(p)}\otimes w_{q+1}^{(p)}+\mathring{R}_{\ell})$ $=\mathord{{\rm div}}\left(\sum_{j,j^{{}^{\prime}},\xi,\xi^{{}^{\prime}}}w_{(\xi)}\otimes w_{(\xi^{{}^{\prime}})}+\mathring{R}_{\ell}\right)$ (4.29)
$=\mathord{{\rm div}}\left(\sum_{j,\xi}c^{-1}_{*}\rho\eta_{k,j}^{2}\gamma_{\xi}^{(j)}\left(\textrm{Id}-c_{*}\rho^{-1}\mathring{R}_{\ell}\right)^{2}B_{\xi}\otimes B_{-\xi}+\mathring{R}_{\ell}\right)$
$\qquad\qquad\qquad+\mathord{{\rm div}}\left(\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}\right)$
$=\mathord{{\rm div}}\left(c^{-1}_{*}\rho(\textrm{Id}-c_{*}\rho^{-1}\mathring{R}_{\ell})+\mathring{R}_{\ell}\right)$
$\qquad\qquad\qquad+\mathord{{\rm div}}\left(\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}\right)$
=$\nabla\left(c_{*}^{-1}\rho\right)+\sum_{j,j^{{}^{\prime}},\xi+\xi^{{}^{\prime}}\not=0}(W_{(\xi)}\otimes W_{(\xi^{{}^{\prime}})})\nabla\left(a_{(\xi)}a_{(\xi^{{}^{\prime}})}\phi_{(\xi)}\phi_{(\xi^{{}^{\prime}})}\right)$
$\qquad\qquad\qquad+\frac{1}{2}\sum_{j,j^{{}^{\prime}},\xi+\xi^{{}^{\prime}}\not=0}a_{(\xi)}a_{(\xi^{{}^{\prime}})}\phi_{(\xi)}\phi_{(\xi^{{}^{\prime}})}\nabla\left(W_{(\xi)}\cdot W_{(\xi^{{}^{\prime}})}\right)$
$=\mathord{{\rm div}}(R_{\textrm{oscillation}})+\nabla p_{\textrm{oscillation}}.$
Note that compared to [BV19], a slight difference in this term is the appearance of $\mathord{{\rm div}}(\rho\mathrm{Id})$, which can be put into the pressure term $p_{\textrm{oscillation}}$. Thus, we have denoted
$p_{\textrm{oscillation}}:=\frac{1}{2}\sum_{j,j^{{}^{\prime}},\xi+\xi^{{}^{\prime}}\not=0}a_{(\xi)}a_{(\xi^{{}^{\prime}})}\phi_{(\xi)}\phi_{(\xi^{{}^{\prime}})}\left(W_{(\xi)}\cdot W_{(\xi^{{}^{\prime}})}\right)+c_{*}^{-1}\rho,$
and
$R_{\textrm{oscillation}}:=\sum_{j,j^{{}^{\prime}},\xi+\xi^{{}^{\prime}}\not=0}\mathcal{R}\left(\left(W_{(\xi)}\otimes W_{(\xi^{{}^{\prime}})}-\frac{W_{(\xi)}\cdot W_{(\xi^{{}^{\prime}})}}{2}\mathrm{Id}\right)\nabla(a_{(\xi)}a_{(\xi^{{}^{\prime}})}\phi_{(\xi)}\phi_{(\xi^{{}^{\prime}})})\right).$
Finally we define the Reynolds stress at the level $q+1$ by
$\mathring{R}_{q+1}=R_{\textrm{transport}}+R_{\textrm{oscillation}}+R_{\textrm{Nash}}+R_{\textrm{corrector}}+R_{\textrm{commutator}}+R_{\textrm{commutator1}},$ (4.30)
and
$p_{q+1}=p_{\ell}-p_{\textrm{oscillation}}-p_{\textrm{corrector}}-p_{\textrm{commutator1}}.$
4.6. Inductive estimates for $\mathring{R}_{q+1}$
In this section, we shall verify (3.8d), (3.8e) and (3.8f) hold ture at the level $q+1$. We first proceed with estimates of $R_{\mathrm{commutator1}}$ in subsection 4.6.1. Then we recall a useful sationary phase lemma in subsection 4.6.2. Finally, the remianing errors are controlled in subsection 4.6.34.6.6
4.6.1. Estimate on $R_{\mathrm{commutator1}}$
First, by the definition of $z_{q}$ we have
$\|z_{q}\|_{C^{0}_{t,x}}\leqslant\lambda_{q+3}^{\frac{\gamma}{4}}.$ (4.31)
Then we use the pathwise estimates (4.26) and (4.31) to obtain the following pathwise bounds
$\|v_{q+1}\mathring{\otimes}(z_{q+1}-z_{\ell})\|_{C^{0}_{t,x}}$ $\leqslant\|v_{q+1}\|_{C^{0}_{t,x}}\|z_{q+1}-z_{\ell}\|_{C^{0}_{t,x}}\leqslant\lambda_{q+3}^{\gamma}\lambda_{q+4}^{\frac{\gamma}{4}}+\lambda_{q+3}^{\frac{5}{4}\gamma}\leqslant\frac{1}{5}\lambda_{q+4}^{\gamma},$
$\|z_{q+1}\mathring{\otimes}z_{q+1}-z_{\ell}\mathring{\otimes}z_{\ell}\|_{C^{0}_{t,x}}$ $\leqslant\|z_{q+1}\|^{2}_{C^{0}_{t,x}}+\|z_{\ell}\|^{2}_{C^{0}_{t,x}}\leqslant\lambda_{q+4}^{\frac{\gamma}{2}}+\lambda_{q+3}^{\frac{\gamma}{2}}\leqslant\frac{1}{5}\lambda_{q+4}^{\gamma},$
$\|\mathcal{R}(z_{q+1}-z_{\ell})\|_{C^{0}_{t,x}}$ $\leqslant\|z_{q+1}-z_{\ell}\|_{C^{0}_{t,x}}\leqslant\lambda_{q+4}^{\frac{\gamma}{4}}+\lambda_{q+3}^{\frac{\gamma}{4}}\leqslant\frac{1}{5}\lambda_{q+4}^{\gamma},$
where we chose $a$ sufficiently large to absorb the constant in the above. Summing over the estimates above, the following pathwise bound holds true for any $t\in\mathbb{R}$
$\|R_{\textrm{commutator1}}\|_{C^{0}_{t,x}}\leqslant 2\|v_{q+1}\mathring{\otimes}(z_{q+1}-z_{\ell})\|_{C^{0}_{t,x}}+\|z_{q+1}$$\mathring{\otimes}z_{q+1}-z_{\ell}\mathring{\otimes}z_{\ell}\|_{C^{0}_{t,x}}$ (4.32)
$+\|\mathcal{R}(z_{q+1}-z_{\ell})\|_{C^{0}_{t,x}}\leqslant\frac{4}{5}\lambda_{q+4}^{\gamma}.$
Next we estimate the $|\!|\!|\cdot|\!|\!|_{C^{0},r}$-norm of $R_{\textrm{com1}}$. Before proceeding, we give an estimate of $z_{q+1}-z_{q}$.
Lemma 4.1.
For any $p\geqslant 1$, we have
$|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},p}\lesssim(\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}+\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{8}})p.$ (4.33)
In particular, for the fixed $r>1$ in Proposition 3.2, we have
$|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},2r}\lesssim(\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}+\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{8}})r,$ (4.34)
where the implicit constants are independent of $q$.
Proof.
By the definition of $z_{q}$, we deduce for any $p\geqslant 1$
$|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},p}$ $=\sup_{t\in\mathbb{R}}\mathbf{E}\left(\sup_{s\in[t,t+1]}\|z_{q+1}(s)-z_{q}(s)\|_{C_{x}^{0}}^{p}\right)^{\frac{1}{p}}$
$\leqslant\sup_{t\in\mathbb{R}}\mathbf{E}\left(\sup_{s\in[t,t+1]}\|\tilde{z}_{q+1}(s)-\tilde{z}_{q}(s)\|_{C_{x}^{0}}^{p}\cdot\chi_{q+1}(\|\tilde{z}_{q+1}(s)\|_{C_{x}^{1}})^{p}\right)^{\frac{1}{p}}$
$+\sup_{t\in\mathbb{R}}\mathbf{E}\left(\sup_{s\in[t,t+1]}\|\tilde{z}_{q}(s)\|_{C_{x}^{0}}^{p}\cdot\left|\chi_{q+1}(\|\tilde{z}_{q+1}(s)\|_{C_{x}^{1}})-\chi_{q}(\|\tilde{z}_{q}(s)\|_{C_{x}^{1}})\right|^{p}\right)^{\frac{1}{p}}$
[…]
$\tilde{z}_{q+1}(t,x)-\tilde{z}_{q}(t,x)=\sum_{\lambda_{q+1}^{\gamma/8}<|k|\leqslant\lambda_{q+2}^{\gamma/8}}e^{ik\cdot x}\hat{z}(t,k),$
where $\hat{z}$ is the Fourier transform of $z$, and $k\in\mathbb{Z}^{3}$. Then using Hölder’s inequality, we deduce for any $t\in{\mathbb{R}}$ and previous $\sigma>0$,
$\|\tilde{z}_{q+1}(t)-\tilde{z}_{q}(t)\|_{L^{\infty}}$ $\leqslant\sum_{\lambda_{q+1}^{\gamma/8}<|k|\leqslant\lambda_{q+2}^{\gamma/8}}|\hat{z}(t,k)|$
$=\sum_{\lambda_{q+1}^{\gamma/8}<|k|\leqslant\lambda_{q+2}^{\gamma/8}}(1+|k|^{2})^{\frac{1}{2}(\frac{3}{2}+\sigma)}|\hat{z}(t,k)|(1+|k|^{2})^{-\frac{1}{2}(\frac{3}{2}+\sigma)}$
$\leqslant\left(\sum_{\lambda_{q+1}^{\gamma/8}<|k|\leqslant\lambda_{q+2}^{\gamma/8}}(1+|k|^{2})^{\frac{3}{2}+\sigma}|\hat{z}(t,k)|^{2}\right)^{\frac{1}{2}}\left(\sum_{\lambda_{q+1}^{\gamma/8}<|k|\leqslant\lambda_{q+2}^{\gamma/8}}(1+|k|^{2})^{-(\frac{3}{2}+\sigma)}\right)^{\frac{1}{2}}$
$\lesssim\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}\|z(t)\|_{H^{\frac{3}{2}+\sigma}}.$
Taking expectation and using (3.7) we obtain the bound
[…]
$\lesssim$ $\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}\sup_{t\in\mathbb{R}}\mathbf{E}\left(\sup_{s\in[t,t+1]}\|z(s)\|_{H^{\frac{3}{2}+\sigma}}^{p}\right)^{\frac{1}{p}}\lesssim\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},2p}\leqslant\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}\sqrt{2p}L.$
[…]
$|\chi_{q+1}$ $(\|\tilde{z}_{q+1}(s)\|_{C_{x}^{1}})-\chi_{q}(\|\tilde{z}_{q}(s)\|_{C_{x}^{1}})|$
$\leqslant$ $\mathbf{1}_{\{\|\tilde{z}_{q}(s)\|_{C_{x}^{1}}\leqslant\lambda_{q+2}^{\gamma},\ \|\tilde{z}_{q+1}(s)\|_{C_{x}^{1}}>\lambda_{q+3}^{\gamma}\}}+\mathbf{1}_{\{\|\tilde{z}_{q}(s)\|_{C_{x}^{1}}>\lambda_{q+2}^{\gamma}\}}$
$\leqslant$ $\mathbf{1}_{\{\|\tilde{z}_{q+1}(s)\|_{C_{x}^{1}}>\lambda_{q+3}^{\gamma}\}}+\mathbf{1}_{\{\|\tilde{z}_{q}(s)\|_{C_{x}^{1}}>\lambda_{q+2}^{\gamma}\}}.$
Using Hölder’s inequality, we obtain
[…]
$+\sup_{t\in\mathbb{R}}\mathbf{E}\left(\sup_{s\in[t,t+1]}\|\tilde{z}_{q}(s)\|_{C_{x}^{0}}^{p}\cdot\mathbf{1}_{\{\|\tilde{z}_{q}(s)\|_{C_{x}^{1}}>\lambda_{q+2}^{\gamma}\}}\right)^{\frac{1}{p}}$
$\leqslant\sup_{t\in\mathbb{R}}\mathbf{E}$ $\left(\sup_{s\in[t,t+1]}\|\tilde{z}_{q}(s)\|_{C_{x}^{0}}^{2p}\right)^{\frac{1}{2p}}\mathbf{P}\left(\|\tilde{z}_{q+1}\|_{C^{0}_{[t,t+1]}C^{1}_{x}}>\lambda_{q+3}^{\gamma}\right)^{\frac{1}{2p}}$
$+\sup_{t\in\mathbb{R}}\mathbf{E}\left(\sup_{s\in[t,t+1]}\|\tilde{z}_{q}(s)\|_{C_{x}^{0}}^{2p}\right)^{\frac{1}{2p}}\mathbf{P}\left(\|\tilde{z}_{q}\|_{C^{0}_{[t,t+1]}C^{1}_{x}}>\lambda_{q+2}^{\gamma}\right)^{\frac{1}{2p}}.$
By Chebyshev’s inequality and (3.7), we have the following bound
$\sup_{t\in\mathbb{R}}\mathbf{P}\left(\|\tilde{z}_{q}\|_{C^{0}_{[t,t+1]}C^{1}_{x}}>\lambda_{q+2}^{\gamma}\right)^{\frac{1}{2p}}$ $\leqslant\lambda_{q+2}^{-\gamma}\sup_{t\in\mathbb{R}}\left(\mathbf{E}\|\tilde{z}_{q}\|^{2p}_{C_{[t,t+1]}^{0}C_{x}^{1}}\right)^{\frac{1}{2p}}$
$\lesssim\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{8}}|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},2p}\leqslant\sqrt{2p}\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{8}}L.$
Similarly as above, we have
$\sup_{t\in\mathbb{R}}\mathbf{P}\left(\|\tilde{z}_{q+1}\|_{C^{0}_{[t,t+1]}C^{1}_{x}}>\lambda_{q+3}^{\gamma}\right)^{\frac{1}{2p}}$ $\leqslant\lambda_{q+3}^{-\gamma}\sup_{t\in\mathbb{R}}\left(\mathbf{E}\|\tilde{z}_{q+1}\|^{2p}_{C_{[t,t+1]}^{0}C_{x}^{1}}\right)^{\frac{1}{2p}}$
$\lesssim\lambda_{q+3}^{-\gamma}\lambda_{q+2}^{\frac{\gamma}{8}}|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},2p}\leqslant\sqrt{2p}\lambda_{q+3}^{-\gamma}\lambda_{q+2}^{\frac{\gamma}{8}}L.$
With the above estimates in hand, using (3.7) again, we obtain
[…]
Finally, combining (4.35) with (4.36) implies
[…]
In particular, taking $p=2r$ impies (4.33). Hence, the proof of Lemma 4.1 is completed. ∎
With Lemma 4.1 in hand, we proceed to estimate $|\!|\!|R_{\textrm{commutator1}}|\!|\!|_{C^{0},r}$, using estimates (4.27), (4.34) and Hölder’s inequality, we have
$|\!|\!|v_{q+1}\mathring{\otimes}(z_{q+1}-z_{q})|\!|\!|_{C^{0},r}\leqslant|\!|\!|v_{q+1}|\!|\!|_{C^{0},p}|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},2r}\lesssim\lambda_{q+1}^{\beta}(\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}+\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{8}}).$
Combining (3.3), (3.7) with (4.34) and using Hölder’s inequality we obtain
$|\!|\!|z_{q+1}\mathring{\otimes}(z_{q+1}-z_{q})|\!|\!|_{C^{0},r}$ $\leqslant|\!|\!|z_{q+1}|\!|\!|_{C^{0},2r}|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},2r}$
$\lesssim|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},2r}|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},2r}\lesssim\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}+\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{8}}.$
Using a standard mollification estimate, we obtain for any $\delta\in(0,\frac{1}{12})$
$\|z_{\ell}(t)-z_{q}(t)\|_{L^{\infty}}\lesssim\ell\|z_{q}\|_{C^{0}_{[t-1,t]}C^{1}_{x}}+\ell^{\frac{1}{2}-\delta}\|z_{q}\|_{C^{1/2-\delta}_{[t-1,t]}C^{0}_{x}}.$
Taking expectation on the both side and using (3.7), we deduce
$|\!|\!|z_{\ell}-z_{q}|\!|\!|_{C^{0},2r}$ $\lesssim\ell\sup_{t\in\mathbb{R}}\left(\mathbf{E}\|z_{q}\|^{2r}_{C_{[t-1,t+1]}^{0}C_{x}^{1}}\right)^{\frac{1}{2r}}+\ell^{\frac{1}{2}-\delta}\sup_{t\in\mathbb{R}}\left(\mathbf{E}\|z_{q}\|^{2r}_{C_{[t-1,t+1]}^{1/2-\delta}C_{x}^{0}}\right)^{\frac{1}{2r}}$ (4.37)
$\lesssim\ell^{\frac{5}{12}}\lambda_{q+1}^{\frac{\gamma}{8}}\sup_{t\in\mathbb{R}}\left(\mathbf{E}\|z\|^{2r}_{C_{[t-1,t+1]}^{1/2-\delta}H^{\frac{3}{2}+\sigma}}\right)^{\frac{1}{2r}}\lesssim\lambda_{q+1}^{\gamma}\lambda_{q}^{-\frac{3}{4}}\sqrt{2r}L.$
Hence, using (4.27), (4.37) and Hölder’s inequality, we deduce
$|\!|\!|v_{q+1}\mathring{\otimes}(z_{\ell}-z_{q})|\!|\!|_{C^{0},r}\leqslant|\!|\!|v_{q+1}|\!|\!|_{C^{0},2r}|\!|\!|z_{\ell}-z_{q}|\!|\!|_{C^{0},2r}\lesssim\lambda_{q+1}^{\beta+\gamma}\lambda_{q}^{-\frac{3}{4}}.$
Similarly, using (3.7) and (4.37) again, we deduce
$|\!|\!|z_{q+1}\mathring{\otimes}(z_{\ell}-z_{q})|\!|\!|_{C^{0},r}$ $\leqslant|\!|\!|z_{q+1}|\!|\!|_{C^{0},2r}|\!|\!|z_{\ell}-z_{q}|\!|\!|_{C^{0},2r}$
$\lesssim|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},2r}|\!|\!|z_{\ell}-z_{q}|\!|\!|_{C^{0},2r}\lesssim\lambda_{q+1}^{\gamma}\lambda_{q}^{-\frac{3}{4}}.$
Summarizing over the estimates above, we obtain
$|\!|\!|R_{\textrm{commutator1}}|\!|\!|_{C^{0},r}$ $\leqslant|\!|\!|v_{q+1}\mathring{\otimes}(z_{\ell}-z_{q})|\!|\!|_{C^{0},r}+|\!|\!|v_{q+1}\mathring{\otimes}(z_{q+1}-z_{q})|\!|\!|_{C^{0},r}+|\!|\!|z_{\ell}-z_{q}|\!|\!|_{C^{0},r}$ (4.38)
$+|\!|\!|z_{q+1}\mathring{\otimes}(z_{\ell}-z_{q})|\!|\!|_{C^{0},r}+|\!|\!|z_{q+1}\mathring{\otimes}(z_{q+1}-z_{q})|\!|\!|_{C^{0},r}+|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},r}$
$\lesssim\lambda_{q+1}^{\beta+\gamma}\lambda_{q}^{-\frac{3}{4}}+\lambda_{q+1}^{\gamma}\lambda_{q}^{-\frac{3}{4}}+\lambda_{q+1}^{\beta}(\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}+\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{8}})$
$\leqslant\frac{1}{48}c_{R}\delta_{q+3}\leqslant\frac{1}{6}\delta_{q+2},$
and
$|\!|\!|\mathring{R}_{\textrm{commutator1}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{48}c_{R}\delta_{q+3}\underline{e},$
where we used $2b^{2}\beta+\beta<\frac{\gamma\sigma}{8}$ to have $\lambda_{q}^{b\beta-\frac{b}{8}\gamma\sigma}<\lambda_{q}^{-2b^{3}\beta}$. We also used $3b\beta<\gamma$, $14b^{3}\beta+2b\gamma<$ 1 to have $\lambda_{q}^{b\beta+b\gamma-b^{2}\gamma}<\lambda_{q}^{-2b^{3}\beta}$ and $\lambda_{q}^{b\gamma+b\beta-\frac{3}{4}}<\lambda_{q}^{-2b^{3}\beta}$ in the above two inequalities and we chose $a$ sufficiently large to absorb the constant.
4.6.2. Stationary phase Lemma
Before we estimate the remaining terms in (4.30) separately to show that the stress $\mathring{R}_{q+1}$ obeys the bound (3.8d), (3.8e) and (3.8f) at level $q+1$, we first introduce the following lemma which makes rigorous the fact that $\mathcal{R}$ obeys the same elliptic regularity estimates as $|\nabla|^{-1}$. We recall the following stationary phase lemma (see for example [BV19, Lemma 5.7] and [DS17, Lemma 2.2]), adapted to our setting.
Lemma 4.2.
Given $\xi\in\mathbb{S}^{2}\cap\mathbb{Q}^{3}$, let $\lambda\xi\in\mathbb{Z}^{3}$ and $\alpha\in(0,1)$. Assume that $a\in C^{m,\alpha}(\mathbb{T}^{3})$ and $\Phi\in C^{m,\alpha}(\mathbb{T}^{3};\mathbb{R}^{3})$ are smooth functions such that the phase function $\Phi$ obeys
$C^{-1}\leqslant|\nabla\Phi|\leqslant C$
on $\mathbb{T}^{3}$, for some constant $C\geqslant 1$. Then, with the inverse divergence operator $\mathcal{R}$ defined in Section 2.1 we have for any $m\in\mathbb{N}$
$\left|\int_{\mathbb{T}^{3}}a(x)e^{i\lambda\xi\cdot\Phi(x)}{\mathord{{\rm d}}}x\right|\lesssim\frac{\|a\|_{C^{m}}+\|a\|_{C^{0}}\|\nabla\Phi\|_{C^{m}}}{\lambda^{m}},$
$\left\|\mathcal{R}\left(a(x)e^{i\lambda\xi\cdot\Phi(x)}\right)\right\|_{C^{\alpha}}\lesssim\frac{\|a\|_{C^{0}}}{\lambda^{1-\alpha}}+\frac{\|a\|_{C^{m,\alpha}}+\|a\|_{C^{0}}\|\nabla\Phi\|_{C^{m,\alpha}}}{{\lambda^{m-\alpha}}},$
where the implicit constant depends on $C,\alpha$ and m (in particular, not on the frequency $\lambda$).
The above lemma is used when estimating the $\|\cdot\|_{C^{0}}$ and $|\!|\!|\cdot|\!|\!|_{C^{0},r}$-norm of the remaining terms in (4.30). Indeed, for a fixed $t\in[k+(j-1)\ell,k+(j+1)\ell]$, in view of the bound (4.8b), we have $\frac{1}{2}\leqslant\|\Phi_{k,j}(\cdot,t)\|_{C^{0}_{x}}\leqslant 2$ on $\mathbb{T}^{3}$. Thus, it follows from Lemma 4.2 that if $a$ is a smooth periodic function, then we have for any $n\in\mathbb{N}$
$\left\|\mathcal{R}(aW_{(\xi)}\circ\Phi_{k,j})\right\|_{C^{0}_{t}C^{\alpha}_{x}}\lesssim\frac{\|a\|_{C^{0}_{t,x}}}{\lambda_{q+1}^{1-\alpha}}+\frac{\|a\|_{C^{0}_{t}C^{n,\alpha}_{x}}+\|a\|_{C^{0}_{t,x}}\|\nabla\Phi_{k,j}\|_{C^{0}_{t}C^{n,\alpha}_{x}}}{{\lambda_{q+1}^{n-\alpha}}}.$ (4.39)
In addition, the same proof that was used in [DS17] to prove Lemma 4.2 gives other useful estimates. Let $\xi\in\Lambda_{j}$ and $\xi^{\prime}\in\Lambda_{j^{\prime}}$ for $|j-j^{\prime}|\leqslant 1$ be such that $\xi+\xi^{\prime}\neq 0$, we have that $|\xi+\xi^{\prime}|\geqslant\bar{c}>0$, for some universal constant $\bar{c}\in(0,1)$. Then for a smooth periodic function $a$, we have the following estimates (see [BV19, (5.37)]), for $n\in\mathbb{N}$
$\left|\int_{\mathbb{T}^{3}}a(x)e^{i\lambda(\xi\cdot\Phi_{k,j}+\xi^{\prime}\cdot\Phi_{k,j^{\prime}})}{\mathord{{\rm d}}}x\right|\lesssim\frac{\|a\|_{C^{0}_{t}C^{n}_{x}}+\|a\|_{C^{0}_{t,x}}(\|\nabla\Phi_{k,j}\|_{C^{0}_{t}C_{x}^{n}}+\|\nabla\Phi_{k,j^{\prime}}\|_{C^{0}_{t}C_{x}^{n}})}{\lambda^{n}},$ (4.40)
and
$\|\mathcal{R}(a(W_{(\xi)}\circ\Phi_{k,j}\otimes$ $W_{(\xi^{\prime})}\circ\Phi_{k,j^{\prime}}))\|_{C^{0}_{t}C_{x}^{\alpha}}$ (4.41)
$\lesssim\frac{\|a\|_{C^{0}_{t,x}}}{\lambda^{1-\alpha}}+\frac{\|a\|_{C^{0}_{t}C_{x}^{n,\alpha}}+\|a\|_{C^{0}_{t,x}}(\|\nabla\Phi_{k,j}\|_{C^{0}_{t}C_{x}^{n,\alpha}}+\|\nabla\Phi_{k,j^{\prime}}\|_{C^{0}_{t}C_{x}^{n,\alpha}})}{{\lambda^{n-\alpha}}}.$
The implicit constant depends only on $\alpha$ and $n$. In the following subsection, we aim to use the above discussion to control the $\|\cdot\|_{C^{0}}$ and $|\!|\!|\cdot|\!|\!|_{C^{0},r}$ norms, respectively, of the transport error, oscillation error, Nash error and corrector error. We first fix an integer $b\geqslant 7$, then choose $\beta$, $\alpha$ sufficiently small such that $14b^{3}\gamma+4b^{3}\beta+2b\alpha<$ 1 and choose $m\in\mathbb{N}$ such that $m>\frac{2b+5}{b-5}$, at last we choose $a$ sufficiently large to absorb the constant.
4.6.3. Estimate on the transport error.
Inspecting the definition of $w_{q+1}^{(p)}$ as (4.16), we notice that the material derivative $\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla$ cannot lead to higher frequency term, since the term $W_{(\xi)}\circ\Phi_{k,j}$ is perfectly transported by $v_{\ell}+z_{\ell}$. Therefore, for $t\in[k,k+1]$, $k\in\mathbb{Z}$ we have
$(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(p)}=\sum_{j,\xi}(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)a_{(\xi)}W_{(\xi)}\circ\Phi_{k,j}.$
Returning to the $C^{N}_{t,x}$-norm estimate of $a_{(\xi)}$ in (4.14) and (4.15) and the definition of $z_{q}$ (3.5), we use chain rule and (3.8a) to obtain that for any $n\in\mathbb{N}_{0}$
$\|(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)a_{(\xi)}\|_{C^{n}_{t,x}}$ $\lesssim\|a_{(\xi)}\|_{C^{n+1}_{t,x}}+\sum_{k=0}^{n}\|v_{\ell}+z_{\ell}\|_{C^{k}_{t,x}}\cdot\|\nabla a_{(\xi)}\|_{C^{n-k}_{t,x}}$ (4.42)
$\lesssim\ell^{-2n-3}\lambda_{q+3}^{(n+2)\gamma}+\sum_{k=0}^{n}\ell^{-k}\lambda_{q+3}^{\gamma}\ell^{-2(n-k+1)-1}\lambda_{q+3}^{(n-k+2)\gamma}$
$\lesssim\ell^{-2n-3}\lambda_{q+3}^{(n+3)\gamma}.$
[…]
$\|R_{\textrm{transport}}\|_{C^{0}_{t,x}}$ $\lesssim\frac{\ell^{-3}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{1-\alpha}}+\frac{\ell^{-2m-5}\lambda_{q+3}^{(m+4)\gamma}+\ell^{-4-m}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{m-\alpha}}$
$\lesssim\frac{\ell^{-3}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{1-\alpha}}\left(1+\frac{\ell^{-2m-2}\lambda_{q+3}^{(m+1)\gamma}}{\lambda_{q+1}^{m-1}}\right)$
$\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3},$
where we used $b\geqslant 7$ and $13b^{3}\gamma+2b^{3}\beta+b\alpha<$ 1 to have $\lambda_{q}^{6+3b^{3}\gamma}<\lambda_{q}^{b-b\alpha-2b^{3}\beta}$, and used $5m+b+5<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-1)b}$, and $a$ was chosen sufficiently large to absorb the constant. Then taking expectation we obtain
$|\!|\!|R_{\textrm{transport}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},$
and
$|\!|\!|R_{\textrm{transport}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$
4.6.4. Estimate on the oscillation error.
Based on previous computations (4.29) in Section 4.5, we can rewrite oscillation error as follows, for $t\in[k,k+1]$, $k\in\mathbb{Z}$
$R_{\textrm{oscillation}}=\sum_{j,j^{{}^{\prime}},\xi+\xi^{{}^{\prime}}\not=0}\mathcal{R}$ $\bigg{[}\left(W_{(\xi)}\circ\Phi_{k,j}\otimes W_{(\xi^{{}^{\prime}})}\circ\Phi_{k,j^{{}^{\prime}}}-\frac{1}{2}(W_{(\xi)}\circ\Phi_{k,j}\cdot W_{(\xi^{{}^{\prime}})}\circ\Phi_{k,j^{{}^{\prime}}})\textrm{Id}\right)$
×$\times\left(\nabla(a_{(\xi)}a_{(\xi^{{}^{\prime}})})+i\lambda_{q+1}a_{(\xi)}a_{(\xi^{{}^{\prime}})}((\nabla\Phi_{k,j}-\textrm{Id})\cdot\xi+(\nabla\Phi_{k,j^{{}^{\prime}}}-\textrm{Id})\cdot\xi^{{}^{\prime}})\right)\bigg{]}.$
Then using (4.14), (4.15) and chain rule, we have that for $n\in\mathbb{N}_{0}$
$\|\nabla(a_{(\xi)}a_{(\xi^{{}^{\prime}})})\|_{C^{n}_{t,x}}$ $\lesssim\sum_{k=0}^{n+1}\|a_{(\xi)}\|_{C^{k}_{t,x}}\|a_{(\xi)}\|_{C^{n-k+1}_{t,x}}$ (4.43)
$\lesssim\sum_{k=0}^{n+1}\ell^{-2k-\frac{1}{2}}\lambda_{q+3}^{(k+1)\gamma}\ell^{-2(n-k+1)-\frac{1}{2}}\lambda_{q+3}^{(n-k+2)\gamma}$
$\lesssim\ell^{-2n-3}\lambda_{q+3}^{(n+3)\gamma}.$
Using (4.8a) and (4.15), we find
$\lambda_{q+1}\|a_{(\xi)}a_{(\xi^{{}^{\prime}})}((\nabla\Phi_{k,j}-\textrm{Id})\cdot\xi+(\nabla\Phi_{k,j^{{}^{\prime}}}-\textrm{Id})\cdot\xi^{{}^{\prime}})\|_{C^{0}_{t,x}}\lesssim\lambda_{q+1}\lambda_{q}^{-\frac{1}{2}}\lambda_{q+3}^{2\gamma}.$ (4.44)
Using (4.8a), (4.14), (4.15), (4.43), (C.7) and chain rule, we obtain for $n\geqslant 1$
$\lambda_{q+1}\|a_{(\xi)}a_{(\xi^{{}^{\prime}})}((\nabla\Phi_{k,j}-\textrm{Id})\cdot\xi+(\nabla\Phi_{k,j^{{}^{\prime}}}-\textrm{Id})\cdot\xi^{{}^{\prime}})\|_{C^{0}_{t}C^{n}_{x}}$ (4.45)
$\lesssim$ $\lambda_{q+1}\sum_{l=0}^{n}\|a_{(\xi)}a_{(\xi^{\prime})}\|_{C^{0}_{t}C^{l}_{x}}\|(\nabla\Phi_{k,j}-\textrm{Id})\cdot\xi+(\nabla\Phi_{k,j^{{}^{\prime}}}-\textrm{Id})\cdot\xi^{{}^{\prime}}\|_{C^{0}_{t}C^{n-l}_{x}}$
$\lesssim$ $\lambda_{q+1}\sum_{l=0}^{n}\ell^{-2l-1}\lambda_{q+3}^{(l+2)\gamma}\ell^{-(n-l)}\lesssim\lambda_{q+1}\ell^{-2n-1}\lambda_{q+3}^{(n+2)\gamma},$
where we used $C^{n}$-norm estimate of $\nabla\Phi_{k,j}$ (C.7) in the second inequality, namely $\|\nabla\Phi_{k,j}(t)\|_{C^{n}_{x}}\leqslant\ell^{-n}$ for $n\geqslant 1$. Combining (4.43), (4.44) with (4.45) and using $C^{n}$-norm estimate of $\nabla\Phi_{k,j}$ (C.7) again, applying (4.41) with $n=m$ and $a$ replaced by $\nabla(a_{(\xi)}a_{(\xi^{{}^{\prime}})})+i\lambda_{q+1}a_{(\xi)}a_{(\xi^{{}^{\prime}})}((\nabla\Phi_{k,j}-\textrm{Id})\cdot\xi+(\nabla\Phi_{k,j^{{}^{\prime}}}-\textrm{Id})\cdot\xi^{{}^{\prime}})$ implies
$\|R_{\textrm{oscillation}}\|_{C^{0}_{t,x}}\lesssim$ $\frac{\ell^{-3}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{1-\alpha}}+\frac{\ell^{-2m-5}\lambda_{q+3}^{(m+4)\gamma}+\ell^{-m-4}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{m-\alpha}}$
$\qquad+\frac{\lambda_{q+1}\lambda_{q}^{-\frac{1}{2}}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}^{1-\alpha}}+\frac{\ell^{-2m-3}\lambda_{q+3}^{(m+3)\gamma}+\ell^{-m-2}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}^{m-1-\alpha}}$
$\lesssim$ $\frac{\ell^{-3}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{1-\alpha}}\left(1+\frac{\ell^{-2m-2}\lambda_{q+3}^{(m+1)\gamma}}{\lambda_{q+1}^{m-2}}\right)+\frac{\lambda_{q+3}^{2\gamma}}{\lambda_{q}^{\frac{1}{2}}\lambda_{q+1}^{-\alpha}}$
$\leqslant$ $\frac{1}{6\cdot 48}c_{R}\delta_{q+3},$
where we used $b\geqslant 7$ and $14b^{3}\gamma+4b^{3}\beta+2b\alpha<$ 1 to have $\lambda_{q}^{6+3b^{3}\gamma}<\lambda_{q}^{b-b\alpha-2b^{3}\beta}$ and $\lambda_{q}^{2b^{3}\gamma}<\lambda_{q}^{\frac{1}{2}-b\alpha-2b^{3}\beta}$, and used $5m+2b+5<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-2)b}$, and we chose $a$ sufficiently large to absorb the constant. Then taking expectation we obtain
$|\!|\!|R_{\textrm{oscillation}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},$
and
$|\!|\!|R_{\textrm{oscillation}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$
4.6.5. Estimate on the Nash error.
Similar to transport error, we write
$w_{q+1}\cdot\nabla(v_{\ell}+z_{\ell})=w_{q+1}^{(p)}\cdot\nabla(v_{\ell}+z_{\ell})+w_{q+1}^{(c)}\cdot\nabla(v_{\ell}+z_{\ell}).$ (4.46)
For the first term on the right hand side, we have for $t\in[k,k+1]$, $k\in\mathbb{Z}$
$w_{q+1}^{(p)}\cdot\nabla(v_{\ell}+z_{\ell})=\sum_{j,\xi}a_{(\xi)}W_{(\xi)}\circ\Phi_{k,j}\cdot\nabla(v_{\ell}+z_{\ell}).$
Similar to (4.42), using (3.8a), (4.14), (4.15), (4.31) and chain rule, we have for $n\in\mathbb{N}_{0}$
$\|a_{(\xi)}\nabla(v_{\ell}+z_{\ell})\|_{C^{n}_{t,x}}\lesssim\ell^{-2n-2}\lambda_{q+3}^{(n+2)\gamma}.$ (4.47)
[…]
$\|\mathcal{R}(w_{q+1}^{(p)}\cdot\nabla(v_{\ell}+z_{\ell}))\|_{C^{0}_{t,x}}\lesssim\frac{\ell^{-2}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}^{1-\alpha}}+\frac{\ell^{-2m-4}\lambda_{q+3}^{(m+3)\gamma}+\ell^{-m-3}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}^{m-\alpha}}.$ (4.48)
The second term on the right hand side of (4.46) is
$w_{q+1}^{(c)}\cdot\nabla(v_{\ell}+z_{\ell})=\sum_{j,\xi}\left(\left(\frac{\nabla a_{(\xi)}}{\lambda_{q+1}}+ia_{(\xi)}(\nabla\Phi_{k,j}-\textrm{Id})\xi\right)\times W_{(\xi)}\circ\Phi_{k,j}\right)\cdot\nabla(v_{\ell}+z_{\ell}).$
Similar as before, we use (3.8a), (4.14), (4.15) and (4.31) to have
$\left\|\frac{\nabla a_{(\xi)}}{\lambda_{q+1}}\cdot\nabla(v_{\ell}+z_{\ell})\right\|_{C^{n}_{t,x}}$ $\lesssim\frac{1}{\lambda_{q+1}}\sum_{k=0}^{n}\|\nabla a_{(\xi)}\|_{C^{k}_{t,x}}\|\nabla(v_{\ell}+z_{\ell})\|_{C^{n-k}_{t,x}}$
$\lesssim\frac{1}{\lambda_{q+1}}\sum_{k=0}^{n}\ell^{-2k-3}\lambda_{q+3}^{(k+2)\gamma}\ell^{-(n-k+1)}\lambda_{q+3}^{\gamma}$
$\lesssim\frac{\ell^{-2n-4}}{\lambda_{q+1}}\lambda_{q+3}^{(n+3)\gamma},$
and we use (4.47) and (C.7) to have
$\|a_{(\xi)}(\nabla\Phi_{k,j}-\textrm{Id})\cdot\nabla(v_{\ell}+z_{\ell})\|_{C^{0}_{t}C^{n}_{x}}$ $\lesssim\sum_{l=0}^{n}\|a_{(\xi)}\nabla(v_{\ell}+z_{\ell})\|_{C^{0}_{t}C^{l}_{x}}\|\nabla\Phi_{k,j}\|_{C^{0}_{t}C_{x}^{n-l}}$
$\lesssim\sum_{l=0}^{n}\ell^{-2l-2}\lambda_{q+3}^{(l+2)\gamma}\ell^{-(n-l)}$
$\lesssim\ell^{-2n-2}\lambda_{q+3}^{(n+2)\gamma}.$
[…]
$\|\mathcal{R}(w_{q+1}^{(c)}\cdot\nabla(v_{\ell}+z_{\ell}))\|_{C^{0}_{t,x}}$ $\lesssim\frac{\ell^{-4}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{2-\alpha}}+\frac{\ell^{-2m-6}\lambda_{q+3}^{(m+4)\gamma}+\ell^{-m-5}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{m+1-\alpha}}$ (4.49)
$\qquad+\frac{\ell^{-2}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}^{1-\alpha}}+\frac{\ell^{-2m-4}\lambda_{q+3}^{(m+3)\gamma}+\ell^{-m-3}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}^{m-\alpha}}.$
Combining (4.48) with (4.49) we obatin
$\|R_{\textrm{Nash}}\|_{C^{0}_{t,x}}$ $\lesssim\|\mathcal{R}(w_{q+1}^{(p)}\cdot\nabla(v_{\ell}+z_{\ell}))\|_{C^{0}_{t,x}}+\|\mathcal{R}(w_{q+1}^{(c)}\cdot\nabla(v_{\ell}+z_{\ell}))\|_{C^{0}_{t,x}}$
$\lesssim\left(\frac{\ell^{-4}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{2-\alpha}}+\frac{\ell^{-2}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}^{1-\alpha}}\right)\left(1+\frac{\ell^{-2m-2}\lambda_{q+3}^{(m+1)\gamma}}{\lambda_{q+1}^{m-1}}\right)\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3},$
where we used $b\geqslant 7$ and $13b^{3}\gamma+3b^{3}\beta+b\alpha<$ 1 to have $\lambda_{q}^{4+2b^{3}\gamma+b\alpha}<\lambda_{q}^{b-2b^{3}\beta}$ and $\lambda_{q}^{8+3b^{3}\gamma+b\alpha}<\lambda_{q}^{2b-2b^{3}\beta}$, and used $5m+b+5<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-1)b}$, and $a$ was chosen sufficiently large to absorb the constant. Then taking expectation, we obtain
$|\!|\!|R_{\textrm{Nash}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},$
and
$|\!|\!|R_{\textrm{Nash}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6\cdot 48}c_{R}\delta_{q+3}\underline{e}.$
4.6.6. Estimate on the corrector error.
The corrector error has two pieces, the transport derivative of $w^{(c)}_{q+1}$ by the flow of $v_{\ell}+z_{\ell}$, and the residual contribution from the nonlinear term, which are easier to estimate due to (4.20) and (4.23). Thus we obtain
$\|w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(c)}\|_{C^{0}_{t,x}}\leqslant\|w_{q+1}^{(c)}\|_{C^{0}_{t,x}}^{2}\leqslant\frac{\lambda_{q}^{8}\lambda_{q+3}^{4\gamma}}{\lambda_{q+1}^{2}}+\lambda_{q+3}^{2\gamma}\lambda_{q}^{-1}\leqslant\frac{1}{3{\cdot}6{\cdot}96}c_{R}\delta_{q+3},$
and
$\|w_{q+1}^{(p)}\mathring{\otimes}w_{q+1}^{(c)}\|_{C^{0}_{t,x}}\lesssim\lambda_{q+3}^{\gamma}\left(\frac{\lambda_{q}^{4}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}}+\lambda_{q+3}^{\gamma}\lambda_{q}^{-\frac{1}{2}}\right)\leqslant\frac{1}{3{\cdot}6{\cdot}96}c_{R}\delta_{q+3},$
where we used $b\geqslant 7$ and $14b^{3}\gamma+4b^{3}\beta<$ 1 to have $\lambda_{q}^{2b^{3}\gamma}<\lambda_{q}^{\frac{1}{2}-2b^{3}\beta}$, $\lambda_{q}^{4+2b^{3}\gamma}<\lambda_{q}^{b-2b^{3}\beta}$ and $\lambda_{q}^{8+4b^{3}\gamma}<\lambda_{q}^{2b-2b^{3}\beta}$, and $a$ was chosen sufficiently large to absorb the constant. Combining the above estimates and taking expectation, we obtain
$|\!|\!|w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(c)}+w_{q+1}^{(p)}\mathring{\otimes}w_{q+1}^{(c)}+w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(p)}|\!|\!|_{C^{0},r}$ $\leqslant\frac{1}{12}\delta_{q+2},$ (4.50)
$|\!|\!|w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(c)}+w_{q+1}^{(p)}\mathring{\otimes}w_{q+1}^{(c)}+w_{q+1}^{(c)}\mathring{\otimes}w_{q+1}^{(p)}|\!|\!|_{L^{1},1}$ $\leqslant\frac{1}{6{\cdot}96}c_{R}\delta_{q+3}\underline{e}.$
The remaining term in corrector error is
$\mathcal{R}\left((\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(c)}\right).$
We use the definition of $w_{q+1}^{(c)}$ to have
$(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(c)}=\sum_{j,\xi}\left((\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)\left(\frac{\nabla a_{(\xi)}}{\lambda_{q+1}}+ia_{(\xi)}(\nabla\Phi_{k,j}-\textrm{Id})\xi\right)\right)\times W_{(\xi)}\circ\Phi_{k,j}.$
Using (3.8a), (4.14) and (4.15), we obtain
$\left\|(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)\frac{\nabla a_{(\xi)}}{\lambda_{q+1}}\right\|_{C^{n}_{t,x}}$ $\lesssim\frac{1}{\lambda_{q+1}}\|a_{(\xi)}\|_{C^{n+2}_{t,x}}+\frac{1}{\lambda_{q+1}}\sum_{k=0}^{n}\|a_{(\xi)}\|_{C^{k+2}_{t,x}}\|v_{\ell}+z_{\ell}\|_{C^{n-k}_{t,x}}$
$\lesssim\frac{1}{\lambda_{q+1}}\ell^{-2n-5}\lambda_{q+3}^{(n+3)\gamma}+\frac{1}{\lambda_{q+1}}\sum_{k=0}^{n}\ell^{-2k-5}\lambda_{q+3}^{(k+3)\gamma}\ell^{-(n-k)}\lambda_{q+3}^{\gamma}$
$\lesssim\frac{\ell^{-2n-5}}{\lambda_{q+1}}\lambda_{q+3}^{(n+4)\gamma}.$
Using (4.42) and (C.7), we obtain
$\left\|(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)(a_{(\xi)}(\nabla\Phi_{k,j}-\textrm{Id}))\right\|_{C^{0}_{t}C^{n}_{x}}$ $\lesssim\sum_{l=0}^{n}\|(\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)a_{(\xi)}\|_{C^{0}_{t}C^{l}_{x}}\|\nabla\Phi_{k,j}\|_{C^{0}_{t}C^{n-l}_{x}}$
$\lesssim\sum_{l=0}^{n}\ell^{-2l-3}\lambda_{q+3}^{(l+3)\gamma}\ell^{-(n-l)}$
$\lesssim\ell^{-2n-3}\lambda_{q+3}^{(n+3)\gamma}.$
[…]
$\left\|\mathcal{R}\left((\partial_{t}+(v_{\ell}+z_{\ell})\cdot\nabla)w_{q+1}^{(c)}\right)\right\|_{C^{0}_{t,x}}\lesssim$ $\frac{\ell^{-5}\lambda_{q+3}^{4\gamma}}{\lambda_{q+1}^{2-\alpha}}+\frac{\ell^{-2m-7}\lambda_{q+3}^{(m+5)\gamma}+\ell^{-m-6}\lambda_{q+3}^{4\gamma}}{\lambda_{q+1}^{m+1-\alpha}}$
$+\frac{\ell^{-3}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{1-\alpha}}+\frac{\ell^{-2m-5}\lambda_{q+3}^{(m+4)\gamma}+\ell^{-m-4}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{m-\alpha}}$
$\lesssim$ $\left(\frac{\ell^{-5}\lambda_{q+3}^{4\gamma}}{\lambda_{q+1}^{2-\alpha}}+\frac{\ell^{-3}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}^{1-\alpha}}\right)\left(1+\frac{\ell^{-2m-2}\lambda_{q+3}^{(m+1)\gamma}}{\lambda_{q+1}^{m-1}}\right)$
$\leqslant$ $\frac{1}{6{\cdot}96}c_{R}\delta_{q+3},$
where we used $b\geqslant 7$ and $14b^{3}\gamma+2b^{3}\beta+b\alpha<$ 1 to have $\lambda_{q}^{10+4b^{3}\gamma}<\lambda_{q}^{2b-\alpha b-2b^{3}\beta}$ and $\lambda_{q}^{6+3b^{3}\gamma}<\lambda_{q}^{b-\alpha b-2b^{3}\beta}$, and used $5m+5+b<bm$ to have $\lambda_{q}^{4m+4+(m+1)b^{3}\gamma}<\lambda_{q}^{(m-1)b}$, and $a$ was chosen sufficiently large to absorb the constant. Taking expetation and combining (4.50), we obatin
$|\!|\!|R_{\textrm{corrector}}|\!|\!|_{C^{0},r}\leqslant\frac{1}{6}\delta_{q+2},$
and
$|\!|\!|R_{\textrm{corrector}}|\!|\!|_{L^{1},1}\leqslant\frac{1}{6{\cdot}48}c_{R}\delta_{q+3}\underline{e}.$
Summarizing over all the estimates we have established above and by (4.3), (4.4) and (4.32), we obtain
$\|\mathring{R}_{q+1}\|_{C^{0}_{t,x}}\leqslant\lambda_{q+4}^{\gamma},\quad|\!|\!|\mathring{R}_{q+1}|\!|\!|_{C^{0},r}\leqslant\delta_{q+2},\quad|\!|\!|\mathring{R}_{q+1}|\!|\!|_{L^{1},1}\leqslant\frac{1}{48}c_{R}\delta_{q+3}\underline{e}.$ (4.51)
Hence, (3.8d), (3.8e) and (3.8f) hold ture at the level $q+1$.
4.7. Inductive estimates for the energy
To conclude the proof of Proposition 3.2, we shall verify (3.9) holds at level $q+1$. To this end, we define
$\delta E(t):=\Big{|}e(t)(1-\delta_{q+2})-\mathbf{E}\|v_{q+1}(t)+{z_{q+1}}(t)\|_{L^{2}}^{2}\Big{|}.$
Proposition 4.3.
It holds for $t\in{\mathbb{R}}$
$\delta E(t)\leqslant$ $\frac{1}{4}\delta_{q+2}e(t).$ (4.52)
Proof.
By the definition of $\zeta_{q}$ in Section 4.3, we find
[…]
which shall be estimated. Let us begin with the bound of the first term on the right hand side of (4.53). Similar as the computation of oscillation error (4.29) in Section 4.5, by (4.13) we have for any $t\in[k,k+1]$, $k\in\mathbb{Z}$
$w_{q+1}^{(p)}\otimes w_{q+1}^{(p)}$
$=\sum_{j,\xi}c^{-1}_{*}\rho\eta_{k,j}^{2}\gamma_{\xi}^{(j)}\left(\textrm{Id}-\frac{c_{*}\mathring{R}_{\ell}}{\rho}\right)^{2}B_{\xi}\otimes B_{-\xi}+\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}$
$=\sum_{j}c^{-1}_{*}\rho\eta_{k,j}^{2}\left(\textrm{Id}-\frac{c_{*}\mathring{R}_{\ell}}{\rho}\right)+\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}$
$=c^{-1}_{*}\rho\textrm{Id}-\mathring{R}_{\ell}+\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}.$
Taking the trace on both sides of the equation and using the fact that $\mathring{R}_{\ell}$ is traceless, we deduce for $t\in[k,k+1]$, $k\in\mathbb{Z}$
$|w_{q+1}^{(p)}|^{2}-3\zeta_{q}=3c_{*}^{-1}\sqrt{\ell^{2}+|\mathring{R}_{\ell}|^{2}}+3(\zeta_{\ell}-\zeta_{q})+\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}\mathrm{Tr}(a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}).$
Hence, we have
$\mathbf{E}\left|\int_{\mathbb{T}^{3}}|w_{q+1}^{(p)}|^{2}-3\zeta_{q}{\mathord{{\rm d}}}x\right|\leqslant$ $3c_{*}^{-1}(2\pi)^{3}\ell+3c_{*}^{-1}\mathbf{E}\|\mathring{R}_{\ell}\|_{L^{1}}+3\cdot(2\pi)^{3}|\zeta_{\ell}-\zeta_{q}|$ (4.54)
$+$ $\mathbf{E}\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}\left|\int_{\mathbb{T}^{3}}\mathrm{Tr}(a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{k,j}\otimes W_{\xi^{\prime}}\circ\Phi_{k,j^{\prime}}){\mathord{{\rm d}}}x\right|.$
In the following, we estimate each term separately, by the choice of parameters, we find
$3c_{*}^{-1}(2\pi)^{3}\ell=3c_{*}^{-1}(2\pi)^{3}\lambda_{q}^{-2}\leqslant\frac{1}{48}\delta_{q+2}e(t),$
which requires $b^{2}\beta<1$ and $a$ sufficiently large to absorb the constant. Using (3.8f) on $\mathring{R}_{q}$, $\operatorname{supp}\varphi_{\ell}\subset[0,\ell]$ and $c_{R}<c_{*}$ we deduce for $t\in{\mathbb{R}}$
$3c_{*}^{-1}\mathbf{E}\|\mathring{R}_{\ell}(t)\|_{L^{1}}\leqslant\frac{1}{16}\delta_{q+2}e(t).$
For the third term in (4.54), we use (3.3), (3.7), (3.8b) and (3.8c) to obtain
$3\cdot(2\pi)^{3}|\zeta_{\ell}-\zeta_{q}|$ $\lesssim\ell\|e^{\prime}\|_{C^{0}_{[t-1,t]}}+\ell\mathbf{E}\|v_{q}\|_{C^{1}_{[t-1,t],x}}(\|v_{q}\|_{C_{[t-1,t],x}^{0}}+{\|z_{q}\|_{C_{[t-1,t],x}^{0}}})$
$\qquad\qquad+\ell^{1/2-\delta}\mathbf{E}\|{z_{q}}\|_{C_{[t-1,t]}^{1/2-\delta}C_{x}^{0}}(\|v_{q}\|_{C_{[t-1,t],x}^{0}}+\|{z_{q}}\|_{C_{[t-1,t],x}^{0}})$
$\lesssim\ell\tilde{e}+\ell\lambda_{q}^{\frac{3}{2}}(|\!|\!|v_{q}|\!|\!|_{C^{0},1}+|\!|\!|z_{q}|\!|\!|_{C^{0},1})+\ell^{\frac{1}{2}-\delta}|\!|\!|z_{q}|\!|\!|_{C^{1/2-\delta}_{t}C^{0}_{x},2}(|\!|\!|v_{q}|\!|\!|_{C^{0},2}+|\!|\!|z_{q}|\!|\!|_{C^{0},2})$
$\lesssim\ell\tilde{e}+\ell\lambda_{q}^{\frac{3}{2}}|\!|\!|v_{q}|\!|\!|_{C^{0},1}+\ell\lambda_{q}^{\frac{3}{2}}|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},1}$
$\qquad\qquad+\ell^{1/2-\delta}|\!|\!|z|\!|\!|_{C^{1/2-\delta}_{t}H^{\frac{3}{2}+\sigma},2}|\!|\!|v_{q}|\!|\!|_{C^{0},2}+\ell^{1/2-\delta}|\!|\!|z|\!|\!|^{2}_{C^{1/2-\delta}_{t}H^{\frac{3}{2}+\sigma},2}$
$\lesssim\ell\tilde{e}+\ell\lambda_{q}^{\frac{3}{2}}\lambda_{q}^{\beta}+\ell^{1/2-\delta}\lambda_{q}^{\beta}\leqslant\frac{1}{48}\delta_{q+2}e(t),$
where we used $14b^{2}\beta+2\beta<$ 1 to have $\lambda_{q}^{-\frac{1}{2}+\beta}<\lambda_{q}^{-2b^{2}\beta}$ and $\lambda_{q}^{-\frac{5}{6}+\beta}<\lambda_{q}^{-2b^{2}\beta}$, and we $a$ was chosen sufficiently large to absorb the constant. For the last term in (4.54), recalling the definition of $\eta_{k,j}$ and $a_{(\xi)}$ in (4.12), we know that for any fixed $t\in[k,k+1]$, the sum with respect to $j,j^{\prime}$ is actually the finite sum of at most two terms. Hence, applying (4.40) with $n=m$ and using the estimates (4.14), (4.15) and (C.7) we obtain
$\mathbf{E}$ $\sum_{j,j^{\prime},\xi+\xi^{\prime}\neq 0}\left|\int_{\mathbb{T}^{3}}\mathrm{Tr}(a_{(\xi)}a_{(\xi^{\prime})}W_{\xi}\circ\Phi_{j}\otimes W_{\xi^{\prime}}\circ\Phi_{j^{\prime}}){\mathord{{\rm d}}}x\right|$
$\lesssim\mathbf{E}\sup_{j.j^{\prime}}\sum_{\xi+\xi^{\prime}\neq 0}\left|\int_{\mathbb{T}^{3}}a_{(\xi)}a_{(\xi^{\prime})}B_{(\xi)}\otimes B_{(\xi^{\prime})}e^{i\lambda_{q+1}(\xi\cdot\Phi_{k,j}+\xi^{\prime}\cdot\Phi_{k,j^{\prime}})}{\mathord{{\rm d}}}x\right|$
$\lesssim\sup_{j,j^{\prime}}\sum_{\xi+\xi^{\prime}\neq 0}\frac{\|a_{(\xi)}a_{(\xi^{\prime})}\|_{C^{0}_{t}C^{m}_{x}}+\|a_{(\xi)}a_{(\xi^{\prime})}\|_{C^{0}_{t,x}}(\|\nabla\Phi_{k,j}\|_{C^{0}_{t}C^{m}_{x}}+\|\nabla\Phi_{k,j^{\prime}}\|_{C^{0}_{t}C^{m}_{x}})}{\lambda_{q+1}^{m}}$
$\lesssim\frac{\ell^{-2m-1}\lambda_{q+3}^{(m+2)\gamma}}{\lambda_{q+1}^{m}}+\frac{\ell^{-m}\lambda_{q+3}^{2\gamma}}{\lambda_{q+1}^{m}}$
$\leqslant\frac{1}{48}\delta_{q+2}e(t),$
where we used $b^{3}\gamma+2b^{2}\beta<1$ and $5m+5<bm$ to have $\lambda_{q}^{4m+2+(m+2)b^{3}\gamma}<\lambda_{q}^{bm-2b^{2}\beta}$, and we chose $a$ sufficiently large to absorb the constant. This completes the bound for (4.54). Going back to (4.53), it remains to control
$\mathbf{E}\left|\int_{\mathbb{T}^{3}}|w_{q+1}^{(c)}|^{2}+2w_{q+1}^{(p)}\cdot w_{q+1}^{(c)}{\mathord{{\rm d}}}x\right|+\mathbf{E}\left|\int_{\mathbb{T}^{3}}|v_{\ell}-v_{q}+z_{q+1}-z_{q}|^{2}{\mathord{{\rm d}}}x\right|$ (4.55)
$\qquad+2\mathbf{E}\left|\int_{\mathbb{T}^{3}}(v_{\ell}-v_{q}+z_{q+1}-z_{q})(v_{q}+z_{q}){\mathord{{\rm d}}}x\right|+2\mathbf{E}\left|\int_{\mathbb{T}^{3}}(v_{\ell}+z_{q+1})\cdot w_{q+1}{\mathord{{\rm d}}}x\right|.$
Using the estimates (4.20) and (4.23), we have
$\mathbf{E}\left|\int_{\mathbb{T}^{3}}|w_{q+1}^{(c)}|^{2}+2w_{q+1}^{(p)}\cdot w_{q+1}^{(c)}{\mathord{{\rm d}}}x\right|\lesssim\mathbf{E}\|w_{q+1}^{(c)}\|_{C^{0}_{t,x}}^{2}+2\mathbf{E}\|w_{q+1}^{(p)}\|_{C^{0}_{t,x}}\|w_{q+1}^{(c)}\|_{C^{0}_{t,x}}$
$\qquad\qquad\lesssim\frac{\lambda_{q}^{8}\lambda_{q+3}^{4\gamma}}{\lambda_{q+1}^{2}}+\lambda_{q+3}^{2\gamma}\lambda_{q}^{-1}+\frac{\lambda_{q}^{4}\lambda_{q+3}^{3\gamma}}{\lambda_{q+1}}+\lambda_{q+3}^{2\gamma}\lambda_{q}^{-\frac{1}{2}}\leqslant\frac{1}{48}\delta_{q+2}e(t),$
where we used $b\geqslant 7$, $14b^{3}\gamma+4b^{2}\beta<$ 1 to have $\lambda_{q}^{2b^{3}\gamma-\frac{1}{2}}<\lambda_{q}^{-2b^{2}\beta}$, $\lambda_{q}^{4+3b^{3}\gamma}<\lambda_{q}^{b-2b^{2}\beta}$ and $\lambda_{q}^{8+4b^{3}\gamma}<\lambda_{q}^{2b-2b^{2}\beta}$, and $a$ was chosen sufficiently large to absorb the constant. Next, we use (3.3), (3.7), (3.8b), (4.6) and (4.33) to obtain
$\mathbf{E}$ $\int_{\mathbb{T}^{3}}|v_{\ell}-v_{q}+z_{q+1}-z_{q}|^{2}{\mathord{{\rm d}}}x+2\mathbf{E}\left|\int_{\mathbb{T}^{3}}(v_{\ell}-v_{q}+z_{q+1}-z_{q})(v_{q}+z_{q}){\mathord{{\rm d}}}x\right|$
$\lesssim\mathbf{E}\left(\|v_{\ell}-v_{q}+z_{q+1}-z_{q}\|_{C_{t,x}^{0}}^{2}\right)+\mathbf{E}\left(\|v_{\ell}-v_{q}+z_{q+1}-z_{q}\|_{C_{t,x}^{0}}\|z_{q}+v_{q}\|_{C_{t,x}^{0}}\right)$
$\lesssim\mathbf{E}\left(\|v_{\ell}-v_{q}\|^{2}_{C_{t,x}^{0}}\right)+|\!|\!|z_{q+1}-z_{q}|\!|\!|^{2}_{C^{0},2}$
$\qquad+\left(|\!|\!|v_{\ell}-v_{q}|\!|\!|_{C^{0},2}+|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{0},2}\right)(|\!|\!|z_{q}|\!|\!|_{C^{0},2}+|\!|\!|v_{q}|\!|\!|_{C^{0},2})$
$\lesssim\ell^{2}\lambda_{q}^{3}+(\lambda_{q+1}^{-\frac{\gamma\sigma}{4}}+\lambda_{q+2}^{-2\gamma}\lambda_{q+1}^{\gamma})+(\ell\lambda_{q}^{\frac{3}{2}}+\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}+\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{4}})(|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},2}+\lambda_{q}^{\beta})$
$\leqslant\lambda_{q}^{-\frac{1}{2}}\lambda_{q+1}^{\beta}+\lambda_{q+1}^{-\frac{\gamma\sigma}{8}+\beta}+\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\beta+\gamma}\leqslant\frac{1}{48}\delta_{q+2}e(t),$
where we used $2b\beta+\beta<\frac{\gamma\sigma}{8}$ to have $\lambda_{q}^{b\beta+2b^{2}\beta}<\lambda_{q}^{\frac{\gamma\sigma}{8}b}$, and used $14b^{2}\beta+2b\beta<$ 1, $3b\beta<\gamma$ to have $\lambda_{q}^{b\beta-\frac{1}{2}}<\lambda_{q}^{-2b^{2}\beta}$ and $\lambda_{q}^{-b^{2}\gamma+b\beta+b\gamma}<\lambda_{q}^{-2b^{2}\beta}$, and $a$ was chosen large enough to absorb the constant. For the last term of (4.55), we recall (cf.(4.19)) that $w_{q+1}$ may be written as the curl of a vector field as
$w_{q+1}=\frac{1}{\lambda_{q+1}}\sum_{j,\xi}\mathrm{curl}(a_{\xi}B_{\xi}e^{i\lambda_{q+1}\xi\cdot\Phi_{k,j}}).$
Integrating by parts and using the estimates (3.3), (3.7), (3.8b) and (4.15), we have
$2\mathbf{E}\left|\int_{\mathbb{T}^{3}}(v_{\ell}+z_{q+1})\cdot w_{q+1}{\mathord{{\rm d}}}x\right|$ $\lesssim\frac{1}{\lambda_{q+1}}\mathbf{E}\left(\sup_{j}\sum_{\xi}\big{\|}a_{\xi}B_{\xi}e^{i\lambda_{q+1}\xi\cdot\Phi_{k,j}}\big{\|}_{C^{0}_{t,x}}\|v_{\ell}+z_{q+1}\|_{C_{t,x}^{1}}\right)$
$\lesssim\frac{1}{\lambda_{q+1}}\sum_{\xi}\mathbf{E}(\|a_{\xi}\|_{C^{0}_{t,x}}\cdot\|v_{\ell}+z_{q+1}\|_{C^{0}_{t}C_{x}^{1}})$
$\lesssim\frac{\lambda_{q+3}^{\gamma}}{\lambda_{q+1}}(\|v_{q}\|_{C^{1}_{[t-1,t+1],x}}+|\!|\!|z_{q+1}|\!|\!|_{C^{1},1})$
$\lesssim\frac{\lambda_{q+3}^{\gamma}}{\lambda_{q+1}}(\lambda_{q}^{\frac{3}{2}}+\lambda_{q+1}^{\frac{\gamma}{8}}|\!|\!|z|\!|\!|_{H^{\frac{3}{2}+\sigma},1})\leqslant\frac{1}{48}\delta_{q+2}e(t),$
where we used $b\geqslant 7$ and $12b^{2}\beta+b^{3}\gamma<$ 1 to have $\lambda_{q}^{b^{3}\gamma+\frac{3}{2}}<\lambda_{q}^{b-2b^{2}\beta}$ and $\lambda_{q}^{b^{3}\gamma+b\gamma}<\lambda_{q}^{b-2b^{2}\beta}$, and $a$ was chosen sufficiently large to absorb the constant. Summing over the above estimates, we obtain
$\Big{|}e(t)(1-\delta_{q+2})-\mathbf{E}\|v_{q+1}(t)+{z_{q+1}}(t)\|_{L^{2}}^{2}\Big{|}\leqslant\frac{1}{4}\delta_{q+2}e(t).$
Hence, the proof of Proposition 3.2 is completed. ∎
5. Stationary solutions to the stochastic Euler equations
We recall that the trajectory space defined in Section 1.1 is $\mathcal{T}=C({\mathbb{R}};C^{\kappa})\times C({\mathbb{R}};C^{\kappa})$ and we take $\kappa=\frac{\vartheta}{2}$ in this section. The corresponding shifts $S_{t}$, $t\in{\mathbb{R}}$, on trajectories are given by
$S_{t}(u,B)(\cdot)=(u(\cdot+t),B(\cdot+t)-B(t)),\quad t\in{\mathbb{R}},\quad(u,B)\in{\mathcal{T}}.$
The notion of stationary solution was introduced in Definition 1.4. Our main result of this section is existence of stationary solutions as limits of ergodic averages of solutions constructed in Section 3. This in particular implies their non-uniqueness.
Theorem 5.1.
Let $u$ be a solution obtained in Theorem 3.3 with $e(t)=K$ for some constant $K\geqslant 6{\cdot 48}{\cdot(2\pi)^{3}}c_{R}^{-1}rL^{2}$ and satisfying (3.12), where $r,L,c_{R}$ are constants as in Theorem 3.3. Then there exists a sequence $T_{n}\to\infty$ and a stationary solution $((\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbf{P}}),\tilde{u},\tilde{B})$ to (1.1) such that
$\frac{1}{T_{n}}\int_{0}^{T_{n}}\mathcal{L}[S_{t}(u,B)]{\mathord{{\rm d}}}t\to\mathcal{L}[\tilde{u},\tilde{B}]$
weakly in the sense of probability measures on $\mathcal{T}$ as $n\to\infty$. Moreover, it holds true that
[…]
Proof.
First, we claim that there exist $\beta^{\prime},\beta^{\prime\prime}>0$ such that for any $N\in\mathbb{N}$
$\sup_{s\in{\mathbb{R}}}\mathbf{E}\|u(\cdot+s)\|^{2r}_{C^{\beta^{\prime}}([-N,N],C^{\beta^{\prime\prime}}_{x})}\lesssim N.$
Indeed, reviewing the proof of Theorem 3.3, and by interpolation we deduce for some $\beta^{\prime}\in(0,\frac{\vartheta}{2})$ and $\beta^{\prime\prime}\in(\frac{\vartheta}{2},\vartheta)$ satisfying $\beta^{\prime}+\beta^{\prime\prime}<\vartheta<{\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot b^{5}}},\frac{1}{21{\cdot}b^{4}}\}}$
$\sum_{q\geqslant 0}|\!|\!|v_{q+1}-v_{q}|\!|\!|_{C^{\beta^{\prime}}_{t}C^{\beta^{\prime\prime}}_{x},2r}$ $\lesssim\sum_{q\geqslant 0}|\!|\!|v_{q+1}-v_{q}|\!|\!|_{C^{\vartheta}_{t,x},2r}$
$\lesssim\sum_{q\geqslant 0}|\!|\!|v_{q+1}-v_{q}|\!|\!|^{1-\vartheta}_{C^{0}_{t,x},2r}|\!|\!|v_{q+1}-v_{q}|\!|\!|^{\vartheta}_{C^{1}_{t,x},2r}$
$\lesssim\sum_{q\geqslant 0}\delta_{q+1}^{\frac{1-\vartheta}{2}}\lambda_{q+1}^{\frac{3}{2}\vartheta}\delta_{q+1}^{\frac{1}{2}\vartheta}\leqslant\sqrt{3r}La^{\frac{3}{2}b\vartheta}+\lambda_{2}^{\beta}\sum_{q\geqslant 1}\lambda_{q+1}^{\frac{3}{2}\vartheta-\beta}<\infty.$
Hence, we conclude that $v=\lim_{q\rightarrow\infty}v_{q}$ exists and lies in $L^{2r}(\Omega,C^{\beta^{\prime}}(\mathbb{R};C^{\beta^{\prime\prime}}))$. Similarly, using (3.3), (3.7), (4.33) and interpolation we deduce for the same $\beta^{\prime}\in(0,\frac{\vartheta}{2})$ and $\beta^{\prime\prime}\in(\frac{\vartheta}{2},\vartheta)$ satisfying $\beta^{\prime}+\beta^{\prime\prime}<\vartheta<{\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot b^{5}}},\frac{1}{21{\cdot}b^{4}}\}}$ and any $p\geqslant 1$
$\sum_{q\geqslant 0}|\!|\!|z_{q+1}-z_{q}|\!|\!|_{C^{\beta^{\prime}}_{t}C^{\beta^{\prime\prime}}_{x},p}$
$\lesssim$ $\sum_{q\geqslant 0}|\!|\!|z_{q+1}-z_{q}|\!|\!|^{(\frac{1}{4}-\beta^{\prime})(1-\beta^{{}^{\prime\prime}})}_{C^{0}_{t,x},p}|\!|\!|z_{q+1}-z_{q}|\!|\!|^{(\frac{1}{4}-\beta^{\prime})\beta^{{}^{\prime\prime}}}_{C^{0}_{t}C^{1}_{x},p}|\!|\!|z_{q+1}-z_{q}|\!|\!|^{\beta^{\prime}(1-\beta^{{}^{\prime\prime}})}_{C^{\frac{1}{4}}_{t}C^{0}_{x},p}|\!|\!|z_{q+1}-z_{q}|\!|\!|^{\beta^{\prime}\beta^{{}^{\prime\prime}}}_{C^{\frac{1}{4}}_{t}C^{1}_{x},p}$
$\lesssim$ $\sum_{q\geqslant 0}(\lambda_{q+1}^{-\frac{\gamma\sigma}{8}}+\lambda_{q+2}^{-\gamma}\lambda_{q+1}^{\frac{\gamma}{8}})^{(\frac{1}{4}-\beta^{\prime})(1-\beta^{{}^{\prime\prime}})}\lambda_{q+2}^{\frac{\gamma}{8}[(\frac{1}{4}-\beta^{\prime})\beta^{\prime\prime}+\beta^{\prime}\beta^{\prime\prime}]}$
$\lesssim$ $\sum_{q\geqslant 0}\lambda_{q+1}^{-\frac{\gamma}{8}{\cdot}\frac{\sigma}{16}}\lambda_{q+2}^{\frac{\gamma}{32}\beta^{\prime\prime}}+\lambda_{q+2}^{-\frac{\gamma}{16}+\frac{\gamma}{32}\beta^{\prime\prime}}\lambda_{q+1}^{\frac{\gamma}{8}}\lesssim\sum_{q\geqslant 0}\lambda_{q+1}^{-\frac{\gamma}{8}{\cdot}\frac{\sigma}{16}+\frac{\gamma}{32}b\beta^{\prime\prime}}+\lambda_{q+1}^{\frac{\gamma}{8}+\frac{\gamma}{32}b\beta^{\prime\prime}-\frac{\gamma}{16}b}<\infty,$
where we used $(\frac{1}{4}-\beta^{\prime})(1-\beta^{{}^{\prime\prime}})>\frac{1}{16}$, $b\geqslant 7$ and $\beta^{\prime\prime}<{\mathord{{\rm min}}}{\{\frac{\sigma}{4b},1\}}$ in the last two inequalities. Hence we obtain $\lim_{q\rightarrow\infty}z_{q}=z$ in $L^{p}(\Omega,C^{\beta^{\prime}}(\mathbb{R};C^{\beta^{\prime\prime}}))$ for any $p\geqslant 1$. Then letting $u=v+z$, we obtain an $(\mathcal{F}_{t})_{t\in\mathbb{R}}$-adapted analytically weak solution $u$ to (1.1) and
$|\!|\!|u|\!|\!|_{C^{\beta^{\prime}}_{t}C^{\beta^{\prime\prime}}_{x},2r}<\infty.$
Hence we obtain
$\sup_{s\in{\mathbb{R}}}\mathbf{E}\|u(\cdot+s)\|^{2r}_{C^{\beta^{\prime}}([-N,N],C^{\beta^{\prime\prime}}_{x})}$ $\leqslant\sup_{s\in{\mathbb{R}}}\sum_{i=-N}^{N-1}\mathbf{E}\|u(\cdot+s)\|^{2r}_{C^{\beta^{\prime}}([i,i+1],C^{\beta^{\prime\prime}}_{x})}$
$\leqslant 2N\sup_{t\in\mathbb{R}}\mathbf{E}\|u\|^{2r}_{C^{\beta^{\prime}}([t,t+1],C^{\beta^{\prime\prime}}_{x})}\lesssim N.$
Now, we define the ergodic averages of the solution $(u,B)$ as the probability measures on the space of trajectories ${\mathcal{T}}$
$\nu_{T}=\frac{1}{T}\int_{0}^{T}\mathcal{L}[S_{t}(u,B)]{\mathord{{\rm d}}}t,\qquad T\geqslant 0.$
For $R_{N}>0$, $N\in\mathbb{N}$, by Arzelà–Ascoli Theorem and the above assertion, we note that the set
$K_{M}:=\cap_{N=M}^{\infty}\bigg{\{}g;\,$ $\|g\|_{C^{\beta^{\prime}}_{[-N,N]}C^{\beta^{\prime\prime}}_{x}}\leqslant R_{N}\bigg{\}}$
is relatively compact in $C({\mathbb{R}};C^{\frac{\vartheta}{2}})$ since $\frac{\vartheta}{2}<\beta^{\prime\prime}$. As a consequence, we deduce that the time shifts $S_{t}u$, $t\in{\mathbb{R}}$, are tight on $C({\mathbb{R}};C^{\frac{\vartheta}{2}})$. Since $S_{t}B$ is a Wiener process for every $t\in{\mathbb{R}}$, the law of $S_{t}B$ does not change with $t\in{\mathbb{R}}$ and is tight. Accordingly, for any $\varepsilon>0$ there is a compact set $\bar{K}_{\varepsilon}$ in ${\mathcal{T}}$ such that
$\sup_{t\in{\mathbb{R}}}{\mathbf{P}}(S_{t}(u,B)\in\bar{K}^{c}_{\varepsilon})<\varepsilon.$
This implies
$\nu_{T}(\bar{K}_{\varepsilon}^{c})=$ $\frac{1}{T}\int_{0}^{T}\mathbf{P}(S_{t}(u,B)\in\bar{K}^{c}_{\varepsilon}){\mathord{{\rm d}}}t<\varepsilon$
and therefore there is a weakly converging subsequence of the probability measures $\nu_{T}$, $T\geqslant 0$. That is, there is a subsequence $T_{n}\to\infty$ and $\nu\in{\mathcal{P}}({\mathcal{T}})$ such that $\nu_{T_{n}}\to\nu$ weakly in ${\mathcal{P}}({\mathcal{T}})$. Define the set
$A$ =$=\Big{\{}(u,B)\in{\mathcal{T}};\ \langle u(t),\psi\rangle+\int_{s}^{t}\langle\mathord{{\rm div}}(u\otimes u),\psi\rangle{\mathord{{\rm d}}}r$
$\qquad\qquad=\langle u(s),\psi\rangle+\langle B(t)-B(s),\psi\rangle,\quad\forall\psi\in C^{\infty}(\mathbb{T}^{3};{\mathbb{R}}^{3}),\ \mathord{{\rm div}}\psi=0,\ t\geqslant s\Big{\}}.$
Since $(u,B)$ in the statement of the theorem satisfies the equation, we have for all $t\in{\mathbb{R}}$
${\mathcal{L}}[S_{t}(u,B)](A)=1.$
Hence, also $\nu_{T_{n}}(A)=1$ for all $n\in\mathbb{N}$. By Jakubowski–Skorokhod representation theorem, there is a probability space $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbf{P}})$ with a sequence of random variables $(\tilde{u}^{n},\tilde{B}^{n})$, $n\in\mathbb{N}$, such that ${\mathcal{L}}[\tilde{u}^{n},\tilde{B}^{n}]={\nu_{T_{n}}}$ and $(\tilde{u}^{n},\tilde{B}^{n})$ satisfy equation (1.1) on ${\mathbb{R}}$. By (3.12) we know
$\tilde{\mathbf{E}}\|\tilde{u}^{n}(t)\|_{L^{2}}^{2}=\frac{1}{T_{n}}\int_{0}^{T_{n}}\mathbf{E}\|{S_{s}u(t)}\|_{L^{2}}^{2}{\mathord{{\rm d}}}s=K.$ (5.2)
By (3.11)
$\sup_{n}\tilde{\mathbf{E}}\|\tilde{u}^{n}(t)\|_{L^{2}}^{2r}=\sup_{n}\frac{1}{T_{n}}\int_{0}^{T_{n}}\mathbf{E}\|S_{s}u(t)\|_{L^{2}}^{2r}{\mathord{{\rm d}}}s<\infty.$ (5.3)
Moreover, there is a random variable $(\tilde{u},\tilde{B})$ having the law ${\mathcal{L}}[\tilde{u},\tilde{B}]=\nu$ so that
$(\tilde{u}^{n},\tilde{B}^{n})\to(\tilde{u},\tilde{B})\qquad\tilde{\mathbf{P}}\text{-a.s. in }\mathcal{T}.$
Thus, we can pass to the limit in the equation to deduce that $\nu$ is a law of a solution on ${\mathbb{R}}$. The shift invariance follows from the same argument as in [BFH20e, Lemma 5.2]. Namely, it holds for $G\in C_{b}({\mathcal{T}})$ and $r\in{\mathbb{R}}$
$\int_{{\mathcal{T}}}G\circ S_{r}(u,B){\mathord{{\rm d}}}\nu(u,B)=\int_{{\mathcal{T}}}G(u,B){\mathord{{\rm d}}}\nu(u,B).$
Finally, (5.1) follows from (5.2) and (5.3). ∎
Using the above result and choosing different $K$, the second claim in Theorem 1.5 follows.
Appendix 1 A Beltrami waves
In this part we recall the Beltrami waves from [BV19, Section 5.4] whcih is adpated to the convex integration scheme in Proposition 3.2. We point out that the construction is entirely deterministic, that is, none of the functions below depends on $\omega$. Let us begin with the definition of Beltrami waves. Given $\xi\in\mathbb{S}^{2}\cap\mathbb{Q}^{3}$, let $A_{\xi}\in\mathbb{S}^{2}\cap\mathbb{Q}^{3}$ obey
$A_{\xi}\cdot\xi=0,\quad A_{-\xi}=A_{\xi}.$
We define the complex vector
$B_{\xi}=\frac{1}{\sqrt{2}}(A_{\xi}+i\xi\times A_{\xi}).$
By construction, the vector $B_{\xi}$ has the properties
$|B_{\xi}|=1,\quad B_{\xi}\cdot\xi=0,\quad i\xi\times B_{\xi}=B_{\xi},\quad B_{-\xi}=\overline{B_{\xi}}.$
This implies that for any $\lambda\in\mathbb{Z}$, such that $\lambda\xi\in\mathbb{Z}^{3}$, the function
$W_{(\xi)}:=W_{\xi,\lambda}(x):=B_{\xi}e^{i\lambda\xi\cdot x}$ (A.1)
is $\mathbb{T}^{3}$ periodic, divergence free, and is an eigenfunction of the $\mathrm{curl}$ operator with eigenvalue $\lambda$. That is, $W_{(\xi)}$ is a complex Beltrami plane wave. The following lemma states a useful property for linear combinations of complex Beltrami plane waves, which can be found in [BV19, Proposition 5.5].
Lemma A.1.
Let $\Lambda$ be a given finite subset of $\mathbb{S}^{2}\cap\mathbb{Q}^{3}$ such that $-\Lambda=\Lambda$, and let $\lambda\in\mathbb{Z}$ be such that $\lambda\Lambda\subset\mathbb{Z}$. Then for any choice of coefficients $a_{\xi}\in\mathbb{C}$ with $\overline{a_{\xi}}=a_{-\xi}$ the vector field
$W(x)=\sum_{\xi\in\Lambda}a_{\xi}B_{\xi}e^{i\lambda\xi\cdot x}$ (A.2)
is a real-valued, divergence-free Beltrami vector field $\mathrm{curl}W=\lambda W$, and thus it is a stationary solution of the Euler equations
$\mathord{{\rm div}}(W\otimes W)=\nabla\frac{|W|^{2}}{2}.$
Furthermore, since $B_{\xi}\otimes B_{-\xi}+B_{-\xi}\otimes B_{\xi}=2Re(B_{\xi}\otimes B_{-\xi})=\mathrm{Id}-\xi\otimes\xi$, we have
[…]
The key point of the construction is that the abundance of Beltrami waves allows to find several such flows $v$ with the property that
$\frac{1}{(2\pi)^{3}}\int_{\mathbb{T}^{3}}v\otimes v(t,x){\mathord{{\rm d}}}x$
equals a prescribed symmetric matrix $R$. Indeed we will need to select these flows so as to depend smoothly on the matrix $R$, at least when $R$ belongs to a neighborhood of the identity matrix. In view of (A.3), such selection is made possible by the following lemma.
Lemma A.2.
(Geometric lemma) There exists a sufficiently small $c_{*}>0$ with the following property. Let $B_{c_{*}(\mathrm{Id})}$ denote the closed ball of symmetric $3\times 3$ matrices, centered at $\mathrm{Id}$, of radius $c_{*}$. Then, there exist pairwise disjoint subsets
$\Lambda_{\alpha}\subset\mathbb{S}^{2}\cap\mathbb{Q}^{3}\qquad\alpha\in\{0,1\},$
and smooth positive functions
$\gamma_{\xi}^{(\alpha)}\in C^{\infty}(B_{c_{*}}(\mathrm{Id}))\qquad\alpha\in\{0,1\},\xi\in\Lambda_{\alpha},$
such that the following hold. For every $\xi\in\Lambda_{\alpha}$ we have $-\xi\in\Lambda_{\alpha}$ and $\gamma_{\xi}^{(\alpha)}=\gamma_{-\xi}^{(\alpha)}$. For each $R\in B_{c_{*}}(\mathrm{Id})$ we have the identity
$R=\frac{1}{2}\sum_{\xi\in\Lambda_{\alpha}}\left(\gamma_{\xi}^{(\alpha)}(R)\right)^{2}(\mathrm{Id}-\xi\otimes\xi).$ (A.4)
We label by $n_{*}$ the smallest natural number such that $n_{*}\Lambda_{\alpha}\subset\mathbb{Z}^{3}$ for all $\alpha\in\{0,1\}$.
[…]
$\sum_{\xi\in\Lambda_{\alpha}}\|\gamma^{(\alpha)}_{\xi}\|_{C^{n}(B_{c_{*}}(\mathrm{Id}))}\leqslant M$ (A.5)
holds for $n$ large enough, $\alpha\in\{0,1\}$ and $\xi\in\Lambda_{\alpha}$. This parameter is universal.
Appendix 2 B Estimates of $a_{(\xi)}$ in the proof of Proposition 3.2
For completeness, we include here the detailed proof of the estimates (4.14) and (4.15). We first aim at estimating the $C^{N}_{t,x}$-norm of $\rho=\sqrt{|\mathring{R}_{\ell}|^{2}+\ell^{2}}+c_{*}\zeta_{\ell}$ for $N\in\mathbb{N}$. To this end, we first apply the chain rule [BDLIS15, Proposition C.1] to the function $\Psi(z)=\sqrt{\ell^{2}+z^{2}}$, since $|D^{m}\Psi(z)|\lesssim\ell^{-m+1}$, for $m\geqslant 1$, we obtain for any $N\in\mathbb{N}$
$\|\rho\|_{C^{N}_{t,x}}$ $\lesssim\left\|\sqrt{\ell^{2}+|\mathring{R}_{\ell}|^{2}}\right\|_{C^{0}_{t,x}}+\|D\Psi\|_{C^{0}}\|\mathring{R}_{\ell}\|_{C^{N}_{t,x}}+\|D\Psi\|_{C^{N-1}}\|\mathring{R}_{\ell}\|^{N}_{C^{1}_{t,x}}+\|\zeta_{\ell}\|_{C^{N}_{t}}$ (B.1)
$\lesssim\ell+\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}+\ell^{-N}\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}+\ell^{-2N+1}\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}^{N}+\ell^{-N}\delta_{q+1}\bar{e},$
where we have used definition of $\zeta_{\ell}$ (4.10) in the last inequality. Let us now estimate the $C^{N}_{t,x}$-norm of $a_{(\xi)}$. By Leibniz rule, we get
$\|a_{(\xi)}\|_{C^{N}_{t,x}}\lesssim\sum_{m=0}^{N}\|\rho^{\frac{1}{2}}\|_{C^{m}_{t,x}}\left\|\gamma_{\xi}\left(\textrm{Id}-\frac{\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{N-m}_{t,x}},$ (B.2)
and estimate each term separately. First, by (4.10)
$\|\rho^{\frac{1}{2}}\|_{C^{0}_{t,x}}\lesssim\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}^{\frac{1}{2}}+\ell^{\frac{1}{2}}+\delta_{q+1}^{\frac{1}{2}}\bar{e}^{\frac{1}{2}},$
and by Lemma A.2
$\left\|\gamma_{\xi}\left(\textrm{Id}-\frac{\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{0}_{t,x}}\lesssim 1.$
Second, applying [BDLIS15, Proposition C.1] to the function $\Psi(z)=z^{1/2},$ $|D^{m}\Psi(z)|\lesssim|z|^{1/2-m},$ for $m=1,\dots,N$, and using (B.1) and $\rho\geqslant\ell$ and $\ell^{-1}\geqslant\bar{e}$ we obtain for $m\geqslant 1$
$\|\rho^{1/2}\|_{C^{m}_{t,x}}\lesssim$ $\|\rho^{1/2}\|_{C^{0}_{t,x}}+\ell^{-1/2}\|\rho\|_{C^{m}_{t,x}}+\ell^{1/2-m}\|\rho\|_{C^{1}_{t,x}}^{m}$ (B.3)
$\lesssim$ $\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}^{\frac{1}{2}}+\ell^{\frac{1}{2}}+\delta_{q+1}^{\frac{1}{2}}\bar{e}^{\frac{1}{2}}+\ell^{\frac{1}{2}-2m}\|\mathring{R}_{q}\|^{m}_{C^{0}_{[t-1,t+1],x}}$
$+\ell^{-\frac{1}{2}-m}\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}+\ell^{\frac{1}{2}-2m}\delta_{q+1}^{m}+\ell^{-\frac{1}{2}-m}\delta_{q+1}.$
For $m\geqslant 1$, using $\delta_{q+1}\leqslant 3rL^{2}$ the above is bounded by
$\ell^{\frac{1}{2}-2m}(3rL^{2}+\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}})^{m}.$
We proceed with a bound for $\left\|\gamma_{\xi}\left(\textrm{Id}-\frac{\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{N-m}_{t,x}}$ for $m=0,\dots,N-1$. Keeping [BDLIS15, Proposition C.1] in mind, we need to estimate
$\left\|\frac{\mathring{R}_{\ell}}{\rho}\right\|_{C^{N-m}_{t,x}}+\left\|\frac{\nabla_{t,x}\mathring{R}_{\ell}}{\rho}\right\|_{C^{0}_{t,x}}^{N-m}+\left\|\frac{\mathring{R}_{\ell}}{\rho^{2}}\right\|_{C^{0}_{t,x}}^{N-m}\|\rho\|_{C^{1}_{t,x}}^{N-m}.$ (B.4)
We use $\rho\geqslant\ell$ to have
$\left\|\frac{\nabla_{t,x}\mathring{R}_{\ell}}{\rho}\right\|_{C^{0}_{t,x}}^{N-m}\lesssim\ell^{-2(N-m)}\|\mathring{R}_{q}\|^{N-m}_{C_{[t-1,t+1],x}^{0}},$ (B.5)
and in view of $|\frac{\mathring{R}_{\ell}}{\rho}|\leqslant 1$
$\left\|\frac{\mathring{R}_{\ell}}{\rho^{2}}\right\|_{C^{0}_{t,x}}^{N-m}\lesssim\left\|\frac{1}{\rho}\right\|_{C^{0}_{t,x}}^{N-m}\lesssim\ell^{-(N-m)},$
and by (B.1) and $\delta_{q+1}\leqslant 3rL^{2}$
$\|\rho\|_{C^{1}_{t,x}}^{N-m}\lesssim\ell^{-(N-m)}\|\mathring{R}_{q}\|^{N-m}_{C_{[t-1,t+1],x}^{0}}+\ell^{-(N-m)},$
hence
$\left\|\frac{\mathring{R}_{\ell}}{\rho^{2}}\right\|_{C^{0}_{t,x}}^{N-m}\|\rho\|_{C^{1}_{t,x}}^{N-m}\lesssim\ell^{-2(N-m)}\|\mathring{R}_{q}\|^{N-m}_{C_{[t-1,t+1],x}^{0}}+\ell^{-2(N-m)}.$ (B.6)
To estimate the first term in (B.4), we write
$\left\|\frac{\mathring{R}_{\ell}}{\rho}\right\|_{C^{N-m}_{t,x}}\lesssim\sum_{k=0}^{N-m}\|\mathring{R}_{\ell}\|_{C^{k}_{t,x}}\left\|\frac{1}{\rho}\right\|_{C^{N-m-k}_{t,x}},$ (B.7)
where for $N-m-k=0$ we have
$\left\|\frac{1}{\rho}\right\|_{C^{0}_{t,x}}\lesssim\ell^{-1},$
and for $k=0,\dots,N-m-1$ using (B.1)
$\left\|\frac{1}{\rho}\right\|_{C^{N-m-k}_{t,x}}$ $\lesssim\left\|\frac{1}{\rho}\right\|_{C^{0}_{t,x}}+\ell^{-2}\|\rho\|_{C^{N-m-k}_{t,x}}+\ell^{-(N-m-k){-1}}\|\rho\|_{C^{1}_{t,x}}^{N-m-k}$
$\lesssim\ell^{-2(N-m-k)-1}+\ell^{-(N-m-k)-2}\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}+\ell^{-2(N-m-k)-1}\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}^{N-m-k}.$
Altogether, we therefore bound (B.7) as
$\left\|\frac{\mathring{R}_{\ell}}{\rho}\right\|_{C^{N-m}_{t,x}}\lesssim\ell^{-2(N-m)-1}$ $\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}+\ell^{-(N-m)-2}\|\mathring{R}_{q}\|^{2}_{C_{[t-1,t+1],x}^{0}}$ (B.8)
$+\ell^{-2(N-m)-1}\|\mathring{R}_{q}\|^{N-m+1}_{C_{[t-1,t+1],x}^{0}}.$
Finally, plugging (B.5), (B.6) and (B.8) into (B.4) leads to
$\left\|\gamma_{\xi}\left(\textrm{Id}-\frac{\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{N-m}_{t,x}}$ $\lesssim\ell^{-2(N-m)-1}\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}+\ell^{-(N-m)-2}\|\mathring{R}_{q}\|^{2}_{C_{[t-1,t+1],x}^{0}}$
$+\ell^{-2(N-m)}\|\mathring{R}_{q}\|^{N-m}_{C_{[t-1,t+1],x}^{0}}+\ell^{-2(N-m)-1}\|\mathring{R}_{q}\|^{N-m+1}_{C_{[t-1,t+1],x}^{0}}+\ell^{-2(N-m)}.$
For $N-m\geqslant 1$ the above is bounded by
$\ell^{-2(N-m)-1}(\|\mathring{R}_{q}\|^{N-m+1}_{C_{[t-1,t+1],x}^{0}}+\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}+1).$
Combining this with the bounds (B.3) above and plugging into (B.2) yields for $N\geqslant 2$
$\|a_{(\xi)}\|_{C^{N}_{t,x}}$ $\lesssim\ell^{-2N-\frac{1}{2}}(1+\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}})^{N+1}\lesssim\ell^{-2N-\frac{1}{2}}\lambda_{q+3}^{(N+1)\gamma},$ (B.9)
where we used inductive assumption (3.8d) for $\mathring{R}_{q}$ in the last inequality. Moreover, for $N=1$, by (B.3) and chain rule, we obtain
$\|a_{(\xi)}\|_{C^{1}_{t,x}}$ $\lesssim\|\rho^{\frac{1}{2}}\|_{C^{1}_{t,x}}+\|\rho^{\frac{1}{2}}\|_{C^{0}_{t,x}}\left\|\gamma_{\xi}\left(\mathrm{Id}-\frac{\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{1}_{t,x}}$ (B.10)
$\lesssim\ell^{-\frac{3}{2}}(\delta_{q+1}+\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}})+(\delta_{q+1}^{\frac{1}{2}}+\ell^{\frac{1}{2}}+\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}^{\frac{1}{2}})\left\|\frac{\mathring{R}_{\ell}}{\rho}\right\|_{C^{1}_{t,x}}.$
Since $\rho\geqslant\ell$ and $|\frac{\mathring{R}_{\ell}}{\rho}|\leqslant 1$, by (B.3) and using chain rule again, we obtain
$\left|\nabla_{t,x}\left(\frac{\mathring{R}_{\ell}}{\rho}\right)\right|\lesssim\frac{|\nabla_{t,x}\mathring{R}_{\ell}|}{\rho}+\frac{|\mathring{R}_{\ell}\cdot\nabla_{t,x}\rho|}{\rho^{2}}\lesssim\ell^{-2}\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}+\ell^{-2}.$ (B.11)
Pluging (B.11) into (B.10), we obtain
$\|a_{(\xi)}\|_{C^{1}_{t,x}}\lesssim\ell^{-2}(1+\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}})^{\frac{3}{2}}\leqslant\ell^{-2}\lambda_{q+3}^{2\gamma}.$ (B.12)
And by the definition of $a_{(\xi)}$, it is easy to get
$\|a_{(\xi)}\|_{C^{0}_{t,x}}$ $\lesssim\|\mathring{R}_{q}\|^{\frac{1}{2}}_{C_{[t-1,t+1],x}^{0}}+\ell^{\frac{1}{2}}+\delta_{q+1}^{\frac{1}{2}}\leqslant\lambda_{q+3}^{\gamma}.$ (B.13)
Hence, (4.14) and (4.15) follow from the estimates (B.9), (B.12) and (B.13).
Appendix 3 C Estimates for transport equantions
We provide a detailed estimate of the solutions to the transport equations in Section 4.3.1. First, we consider the following transport equation on $[t_{0},T],t_{0}\geqslant 0$:
$(\partial_{t}+v\cdot\nabla)f$ $=g,$ (C.1)
$f(t_{0},x)$ $=f_{0}.$
We recall the following results from [BDLIS15, Proposition D.1, (133), (134)]: {equ} ∥f(t)∥_C_x^1⩽∥f_0∥_C_x^1e^(t-t_0)∥v∥_C^0_[t_0,t]C^1_x+∫_t_0^te^(t-τ)∥v∥_C_[t_0,t]^0C_x^1∥g(τ)∥_C_x^1 dτ, and more generally, for any $N\geqslant 2$, there exists a constant $C=C_{N}$ such that
$\|f(t)\|_{C_{x}^{N}}\leqslant$ $\left(\|f_{0}\|_{C_{x}^{N}}+C(t-t_{0})\|{v}\|_{C^{0}_{[t_{0},t]}C^{N}_{x}}\|f_{0}\|_{C_{x}^{1}}\right)e^{C(t-t_{0})\|{v}\|_{C^{0}_{[t_{0},t]}C^{1}_{x}}}$ (C.2)
$+\int_{t_{0}}^{t}e^{C(t-\tau)\|{v}\|_{C^{0}_{[t_{0},t]}C^{1}_{x}}}\left(\|g(\tau)\|_{C_{x}^{N}}+C(t-\tau)\|{v}\|_{C^{0}_{[t_{0},t]}C^{N}_{x}}\|g(\tau)\|_{C_{x}^{1}}\right){\mathord{{\rm d}}}\tau.$
Consider the following case:
$(\partial_{t}+v\cdot\nabla)\Phi$ $=0,$
$\Phi(t_{0},x)$ $=x.$
We aim at estimating the $C_{x}^{0}$-norm of $\nabla\Phi(t)-\textrm{Id}$. First, let $\Psi(s,x)=\Phi(s,x)-x$, then $\Psi$ statifies the following equation:
$(\partial_{t}+v\cdot\nabla)\Psi=-v,$
$\Psi(t_{0},x)=0.$
We then apply (C) to conclude
$\|\nabla\Phi(t)-\textrm{Id}\|_{C_{x}^{0}}=\|\Psi(t)\|_{C_{x}^{1}}\leqslant\int_{t_{0}}^{t}e^{(t-\tau)\|{v}\|_{C_{[t_{0},t]}^{0}C_{x}^{1}}}\|{v}\|_{C^{0}_{[t_{0},t]}C_{x}^{1}}{\mathord{{\rm d}}}\tau=e^{(t-t_{0})\|{v}\|_{C_{[t_{0},t]}^{0}C_{x}^{1}}}-1.$ (C.3)
Now we use the above result to estimate the solution $\Phi_{k,j}$ of (4.7), $j\in\{0,1,...,\lceil\ell^{-1}\rceil\}$. Using (3.8c) and the definition of $z_{q}$ in (3.5), we obtain for any $k\in\mathbb{Z}$
$\ell\|{v_{\ell}+z_{\ell}\|_{C^{0}_{[k,k+1]}C_{x}^{1}}}\leqslant\ell(\lambda_{q}^{\frac{3}{2}}\delta_{q}^{\frac{1}{2}}+\lambda_{q+3}^{\frac{\gamma}{4}})\lesssim\lambda_{q}^{-\frac{1}{2}}\ll 1,$ (C.4)
where we used $b^{3}\gamma<1$ in the second inequality. Since $e^{x}-1\leqslant 2x$ for $x\in[0,1]$ and combining (C.3) with (C.4), we obtain
$\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)-\textrm{Id}\|_{C_{x}^{0}}$ $\leqslant e^{2\ell\|v_{\ell}+z_{\ell}\|_{C^{0}_{[k,k+1]}C_{x}^{1}}}-1$ (C.5)
$\lesssim\ell\|{v_{\ell}+z_{\ell}\|_{C^{0}_{[k,k+1]}C_{x}^{1}}}\lesssim\lambda_{q}^{-\frac{1}{2}}\ll 1,$
hence (4.8a) holds ture. Then, (4.8b) follows from (C.5), we have
$\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)\|_{C_{x}^{0}}\leqslant\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)-\mathrm{Id}\|_{C_{x}^{0}}+1\leqslant 2,$ (C.6)
$\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)\|_{C_{x}^{0}}\geqslant 1-\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)-\mathrm{Id}\|_{C_{x}^{0}}\geqslant\frac{1}{2}.$
Similarly, we use (C.2) and (C.4) again, we obtain for any $n\geqslant 1$
$\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla\Phi_{k,j}(t)\|_{C_{x}^{n}}$ $\lesssim\ell\|{v_{\ell}+z_{\ell}\|_{C^{0}_{[k,k+1]}C_{x}^{n+1}}}e^{2\ell\|{v_{\ell}+z_{\ell}\|_{C^{0}_{[k,k+1]}C_{x}^{1}}}}$ (C.7)
$\lesssim\ell^{-(n-1)}\|v_{\ell}+z_{\ell}\|_{C^{0}_{[k,k+1]}C_{x}^{1}}\lesssim\ell^{-n+\frac{1}{4}},$
where we used $b^{3}\gamma<1$ in the last inequality. Using (3.8a), (4.7), (4.31) and (C.6), we obtain
$\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\partial_{t}\Phi_{k,j}\|_{C_{x}^{0}}\leqslant\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|(v_{\ell}+z_{\ell})\cdot\nabla\Phi_{k,j}\|_{C_{x}^{0}}\leqslant\lambda_{q+3}^{\frac{\gamma}{2}}.$ (C.8)
Differentiating both sides of (4.7) and using (3.8a), (3.8c), (4.31), (C.6) and (C.7), we can also obtain
$\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\partial_{t}\nabla\Phi_{k,j}\|_{C_{x}^{0}}\lesssim$ $\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|\nabla(v_{\ell}+z_{\ell})\cdot\nabla\Phi_{k,j}\|_{C_{x}^{0}}$ (C.9)
$+\sup_{t\in[k+(j-1)\ell,k+(j+1)\ell]}\|(v_{\ell}+z_{\ell})\cdot\nabla^{2}\Phi_{k,j}\|_{C_{x}^{0}}$
$\lesssim$ $\lambda_{q}^{\frac{3}{2}}\delta_{q}^{\frac{1}{2}}+\lambda_{q+3}^{\frac{\gamma}{4}}+(\lambda_{q+3}^{\frac{\gamma}{4}}+\lambda_{q+2}^{\gamma})\ell^{-\frac{3}{4}}\leqslant\lambda_{q}^{\frac{3}{2}}\lambda_{q+3}^{\gamma}.$
The estimates (C.7)–(C.9) are used in Section 4.3 and Section 4.6 to estimate the $C^{1}_{t,x}$-norm of $w_{q+1}$ and the $C^{0}_{t,x}$-norm of $\mathring{R}_{q+1}$.